### v3

# Feng-Yu Wanga), Chenggui Yuanb), Xiao-Yu Zhaoa)

Abstract
We consider stochastic differential equations on $\mathbb{R}^{d}$ with coefficients depending on the path and distribution for the whole history. Under a local integrability condition on the time-spatial singular drift, the well-posedness and Lipschitz continuity in initial values are proved, which is new even in the distribution independent case. Moreover, under a monotone condition, the asymptotic log-Harnack inequality is established, which extends the corresponding result of [5] derived in the distribution independent case.
Keywords: Path-distribution dependent SDEs, well-posedness, asymptotic log-Harnack inequality, gradient estimate.
1 Introduction
The dimension-free Harnack inequality with power was first introduced in [11] to study the log-Sobolev inequality on Riemannian manifolds, and has been extensively developed and applied to derive various regularity estimates for SDEs (stochastic differential equations), SPDEs (stochastic partial differential equations), path dependent SDEs, and distribution-dependent SDEs, see [9, 10, 13] and references therein. As a limit version when the power goes to infinite, the log-Harnack inequality has been introduced in [12] to characterize the curvature lower bounded and entropy-cost estimates, and has been extended to metric measure spaces [1, 8]. When the noise of a stochastic system is weak such that the log-Harnack inequality is not available, the asymptotic log-Harnack inequality has been studied in [5, 16] for path-dependent SDEs and SPDEs, which in particular implies an asymptotic gradient estimate.
In this paper, we study the well-posedness and asymptotic log-Harnack inequality for path-distribution dependent SDEs with infinite memory.
[…]
To describe the path dependence with exponential decay memory, let $\mathscr{C}:=C((-\infty,0];\mathbb{R}^{d})$ and for $\tau>0$, set
$\mathscr{C}_{\tau}=\left\{\xi\in\mathscr{C}:\|\xi\|_{\tau}:=\sup_{s\in(-\infty,0]}(\text{\rm{e}}^{\tau s}|\xi(s)|)<\infty\right\}.$ (1.1)
It is well known that $(C((-\infty,0];\mathbb{R}^{d}),\|\cdot\|_{\infty})$ is complete but not separable, so is $(\mathscr{C}_{\tau},\|\cdot\|_{\tau})$ due to the isometric
$\mathscr{C}_{\tau}\ni\xi:=(\xi_{s})_{s\in(-\infty,0]}\mapsto\text{\rm{e}}^{\tau\cdot}\xi:=(\text{\rm{e}}^{\tau s}\xi_{s})_{s\in(-\infty,0]}\in C((-\infty,0];\mathbb{R}^{d}).$
Let $\mathscr{P}$ and $\mathscr{P}(\mathscr{C}_{\tau})$ be the set of all probability measures on $(\mathbb{R}^{d},\mathscr{B}(\mathbb{R}^{d}))$ and $(\mathscr{C}_{\tau},\mathscr{B}(\mathscr{C}_{\tau}))$, respectively, equipped with the weak topology. Let $\mathscr{B}_{b}(\mathscr{C}_{\tau})$ be the class of bounded measurable functions on $\mathscr{C}_{\tau}$, and $\mathscr{B}_{b}^{+}(\mathscr{C}_{\tau})$ the set of strictly positive functions in $\mathscr{B}_{b}(\mathscr{C}_{\tau})$.
Let $(W(t))_{t\geq 0}$ be a $d$-dimensional Brownian motion defined on a complete filtered probability space $(\Omega,\mathscr{F},(\mathscr{F}_{t})_{t\geq 0},\mathbb{P})$. For an $\mathscr{F}_{0}$-measurable random variable $X_{0}:=((-\infty,0]\ni r\mapsto X(r))$ on $\mathscr{C}_{\tau}$, we consider the following path-distribution dependent SDE with infinite memory:
$\text{\rm{d}}X(t)=b(t,X_{t},\mathscr{L}_{X_{t}})\text{\rm{d}}t+\sigma(t,X_{t})\mathrm{d}W(t),\quad t\geq 0,$ (1.2)
where for each fixed $t\geq 0$, $X_{t}(\cdot)\in\mathscr{C}_{\tau}$ is defined by
$X_{t}(r):=X(t+r),\ \ r\in(-\infty,0],$
which is called the segment process of $X(t)$, $\mathscr{L}_{X_{t}}\in\mathscr{P}(\mathscr{C}_{\tau})$ is the distribution of $X_{t}$, and
$b:\mathbb{R}^{+}\times\mathscr{C}_{\tau}\times\mathscr{P}(\mathscr{C}_{\tau})\rightarrow\mathbb{R}^{d},\ \ \ \sigma:\mathbb{R}^{+}\times\mathscr{C}_{\tau}\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{d}$
are measurable. When different probability spaces are concerned, we use $\mathscr{L}_{X_{t}|\mathbb{P}}$ in place of $\mathscr{L}_{X_{t}}$ to emphasize the underline probability.
For any constant $k\geq 0$, let
$\mathscr{P}_{k}(\mathscr{C}_{\tau}):=\big{\{}\mu\in\mathscr{P}(\mathscr{C}_{\tau}):\|\mu\|_{k}:=\mu(\|\cdot\|_{\tau}^{k})^{\frac{1}{k}}<\infty\big{\}},$
where for $k=0$ we set $\mu(\|\cdot\|_{\tau}^{k})^{\frac{1}{k}}=1$ such that $\mathscr{P}_{0}(\mathscr{C}_{\tau})=\mathscr{P}(\mathscr{C}_{\tau}).$ When $k>0$, $\mathscr{P}_{k}(\mathscr{C}_{\tau})$ is a complete (but not separable) space under the $L^{k}$-Wasserstein distance,
$\mathbb{W}_{k}(\mu,\nu):=\inf_{\pi\in C(\mu,\nu)}\bigg{(}\int_{\mathscr{C}_{\tau}\times\mathscr{C}_{\tau}}\|\xi-\eta\|_{\tau}^{k}\,\pi(\mathrm{d}\xi,\mathrm{d}\eta)\bigg{)}^{\frac{1}{1\vee k}},\ \ \mu,\nu\in\mathscr{P}_{k}(\mathscr{C}_{\tau}),$
where $C(\mu,\nu)$ is the set of all coupling of $\mu$ and $\nu$. Moreover, for any $k\geq 0$, $\mathscr{P}_{k}(\mathscr{C}_{\tau})$ is also a complete (but not separable) metric space under the weighted variation distance
$\|\mu-\nu\|_{k,var}:=\sup_{f\in\mathscr{B}_{b}(\mathscr{C}_{\tau}),|f|\leq 1+\|\cdot\|_{\tau}^{k}}|\mu(f)-\nu(f)|=|\mu-\nu|(1+\|\cdot\|_{\tau}^{k}),$
where $|\mu-\nu|$ is the total variation of $\mu-\nu$. According to [10, Remark 3.2.1], for any $k>0$, there exists a constant $c>0$ such that
$\|\mu-\nu\|_{var}+\mathbb{W}_{k}(\mu,\nu)^{1\vee k}\leq c\|\mu-\nu\|_{k,var},\ \ \mu,\nu\in\mathscr{P}_{k}(\mathscr{C}_{\tau}).$ (1.3)
Denote $\mathbb{W}_{k,var}=\mathbb{W}_{k}+\|\cdot\|_{k,var}$ for simplicity.
We will solve (1.2) for distribution $\mathscr{L}_{X_{t}}$ belonging to a subclass $\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})\subset\mathscr{P}_{k}(\mathscr{C}_{\tau})$ such that $\mathscr{L}_{X_{t}}$ is weakly continuous in $t\geq 0$.
Definition 1.1.
  1. (1)
    An adapted continuous process $(X_{t})_{t\geq 0}$ on $\mathscr{C}_{\tau}$ is called a segment solution of (1.2) with initial value $X_{0}$, if $X_{0}$ is an $\mathscr{F}_{0}$-measurable random variable on $\mathscr{C}_{\tau}$, and $(X(t):=X_{t}(0))_{t\geq 0}$ satisfies
    $\int_{0}^{t}\mathbb{E}\big{[}|b(r,X_{r},\mathscr{L}_{X_{r}})|+\|\sigma(r,X_{r})\|^{2}\big{|}\mathscr{F}_{0}\big{]}\,\mathrm{d}r<\infty,\ \ t\geq 0,$
    and $\mathbb{P}$-a.s.
    $X(t)=X(0)+\int_{0}^{t}b(r,X_{r},\mathscr{L}_{X_{r}})\,\mathrm{d}r+\int_{0}^{t}\sigma(r,X_{r})\,\mathrm{d}W(r),\ \ t\geq 0.$
    In this case $(X(t))_{t\geq 0}$ is called the (strong) solution. The SDE (1.2) is called strongly well-posed for distributions in $\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})$, if for any $\mathscr{F}_{0}$-measurable $X_{0}$ with $\mathscr{L}_{X_{0}}\in\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})$, it has a unique segment solution with $\mathscr{L}_{X_{t}}\in\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})$ for $t\geq 0$.
  2. (2)
    A couple $(X_{t},W(t))_{t\geq 0}$ is called a weak segment solution of (1.2) with initial distribution $\mu\in\tilde{\mathscr{P}_{k}}(\mathscr{C}_{\tau}),$ if there exists a probability space under which $W(t)$ is $d$-dimensional Brownian motion and $\mathscr{L}_{X_{0}}=\mu$, such that $(X(t):=X_{t}(0))_{t\geq 0}$ solves the SDE (1.2). We call (1.2) weakly unique, if for any two weak segment solutions $(X_{t}^{i},W^{i}(t))_{t\geq 0}$ under probabilities $\mathbb{P}^{i}$ with common initial distribution, we have $\mathscr{L}_{X_{t}^{1}|\mathbb{P}^{1}}=\mathscr{L}_{X_{t}^{2}|\mathbb{P}^{2}}$ for all $t\geq 0$. We call (1.2) weakly well-posed for distributions in $\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})$, if for any initial distribution it has a unique weak segment solution.
  3. (3)
    The SDE (1.2) is called well-posed for distributions in $\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})$, if it is both strongly and weakly well-posed for distributions in $\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})$. In this case, for any $\xi\in\mathscr{C}_{\tau}$ such that $\mathscr{L}_{\xi}\in\tilde{\mathscr{P}}_{k}(\mathscr{C}_{\tau})$, let
    $P_{t}^{*}\gamma=\mathscr{L}_{X_{t}^{\xi}},\ \ \gamma=\mathscr{L}_{\xi}$
    and denote
    $P_{t}f(\gamma):=\mathbb{E}[f(X_{t}^{\xi})]=\int_{\mathscr{C}_{\tau}}f\,\text{\rm{d}}\left\{P_{t}^{*}\gamma\right\},\ \ \gamma=\mathscr{L}_{\xi},\,t\geq 0,\,f\in\mathscr{B}_{b}(\mathscr{C}_{\tau}),$
    where $(X_{t}^{\xi})_{t\geq 0}$ is the unique solution to (1.2) with $X_{0}^{\xi}=\xi$.
To characterize the singularity of coefficients in time–space variables, we recall some functional spaces introduced in [14]. For any $p\geq 1$, $L^{p}(\mathbb{R}^{d})$ is the class of measurable functions $f$ on $\mathbb{R}^{d}$ such that
$\|f\|_{L^{p}(\mathbb{R}^{d})}:=\bigg{(}\int_{\mathbb{R}^{d}}|f(x)|^{p}\,\mathrm{d}x\bigg{)}^{\frac{1}{p}}<\infty.$
For any $p,q\geq 1$ and $0\leq s<t$, let $\tilde{L}_{q}^{p}(s,t)$ denote the class of measurable functions $f$ on $[s,t]\times\mathbb{R}^{d}$ such that
$\|f\|_{\tilde{L}_{q}^{p}(s,t)}:=\sup_{z\in\mathbb{R}^{d}}\bigg{(}\int_{s}^{t}\|1_{B(z,1)}f_{r}\|_{L^{p}(\mathbb{R}^{d})}^{q}\,\mathrm{d}r\bigg{)}^{\frac{1}{q}}<\infty,$
where $B(z,1):=\{x\in\mathbb{R}^{d}:|x-z|\leq 1\}$.
When $s=0,$ we simply denote
$\tilde{L}_{q}^{p}(t)=\tilde{L}_{q}^{p}(0,t),\ \ t>0.$
We will take $(p,q)$ from the class
$\mathscr{K}:=\Big{\{}(p,q):p,q\in({2},\infty),\ \frac{d}{p}+\frac{2}{q}<1\Big{\}}.$
2 The Singular Case: Well-Posedness and Lipschitz Continuity in Initial Value
In this section, we let $k\geq 0$ and consider (1.2) with singular drifts and $\sigma(t,\xi)=\sigma(t,\xi(0))$.
2.1 Path Dependent SDEs with Infinite Memory
In this part, we consider the following path dependent SDE with infinite memory on $\mathbb{R}^{d}$:
$\text{\rm{d}}X(t)=b(t,X_{t})\mathrm{d}t+\sigma(t,X(t))\mathrm{d}W(t),\ \ t>0,\ \ X_{0}=\xi\in\mathscr{C}_{\tau}.$ (2.1)
To ensure the existence and uniqueness of solutions to (2.1), we decompose $b$ as
$b(t,\xi)=b^{(0)}(t,\xi(0))+b^{(1)}(t,\xi),\ \ t\geq 0,\ \xi\in\mathscr{C}_{\tau}$
and make the following assumptions on $b^{(0)},b^{(1)}$ and $\sigma$. For any $\xi\in\mathscr{C}_{\tau}$, let $\xi^{0}\in\mathscr{C}_{\tau}$ be defined as
$\xi^{0}(r)=\xi(0),\ \ r\leq 0.$
  1. ($A_{1}$)
    $a:=\sigma\sigma^{*}$ is invertible with $\|a\|_{\infty}+\|(a)^{-1}\|_{\infty}<K$ for some constant $K>0$ and
    $\lim_{\varepsilon\downarrow 0}\sup_{|x-y|\leq\varepsilon,t\in[0,T]}{\|a_{t}(x)-a_{t}(y)\|}=0,\quad T\in(0,\infty),\ x,y\in\mathbb{R}^{d}.$
  2. ($A_{2}$)
    There exist constants $\{(p_{i},q_{i})\}_{0\leq i\leq l}\in\mathscr{K}$ with $l\geq 1$, $p_{i}>2$, and functions $0\leq f_{i}\in\cap_{n\in\mathbb{N}}\tilde{L}^{p_{i}}_{q_{i}}(n)$, $0\leq i\leq l$, such that
    $|b^{(0)}|\leq f_{0},\ \ \|\nabla\sigma\|\leq\sum_{i=0}^{l}{f_{i}}.$
  3. $(A_{3})$
    For every $n>0$, there exists a constant $K_{n}>0$ such that
    $|b^{(1)}(t,\xi)-b^{(1)}(t,\eta)|\leq K_{n}\|\xi-\eta\|_{\tau},\ \ {t\geq 0},\,\|\xi\|_{\tau},\|\eta\|_{\tau}\leq n.$ (2.2)
    Moreover, $\sup_{t\geq 0}|b^{(1)}(t,{\bf 0})|<\infty$, where ${\bf 0}\in\mathscr{C}_{\tau}$ with ${\bf 0}(r)=0,r\leq 0$, and there exists a constant $K>0$ such that
    $|b^{(1)}(t,\xi)-b^{(1)}(t,\xi^{0})|\leq K(1+\|\xi\|_{\tau}),\ \ {t\geq 0},\,\xi\in\mathscr{C}_{\tau}.$
Theorem 2.1.
Assume $(A_{1})$$(A_{3})$. Then (2.1) is well-posed for any initial value in $\mathscr{C}_{\tau}$, and for any constants $k,T>0$, there exists a constant $c>0$, such that
$\mathbb{E}\left[\sup_{t\in[0,T]}\left(1+\|X_{t}\|_{\tau}^{k}\right)\bigg{|}X_{0}\right]\leq c\left(1+\|X_{0}\|^{k}_{\tau}\right).$ (2.3)
Let $X_{t}^{\xi}$ be the segment solution with $X_{0}^{\xi}=\xi$. To ensure the Lipschitz continuity of $X_{t}^{\xi}$ in $\xi$, we strengthen $(A_{3})$ to
  • $(A_{3}^{\prime})$
    $\sup_{t\geq 0}|b^{(1)}(t,{\bf 0})|<\infty$, and there exist constants $K>0$ and $\alpha\in[0,1]$ such that
    $|b^{(1)}(t,\xi)-b^{(1)}(t,\eta)|\leq K\|\xi-\eta\|_{\tau},$
    $|b^{(1)}(t,\xi)-b^{(1)}(t,\xi^{0})|\leq K\left(1+\|\xi\|_{\tau}^{\alpha}\right),\ \ t\geq 0,\ \xi,\eta\in\mathscr{C}_{\tau}.$
Theorem 2.2.
Assume $(A_{1})$, $(A_{2})$ and $(A_{3}^{\prime})$. Then for any constants $\varepsilon\in(0,1)$ and $k,T\geq 1$, there exists a constant $c>0$ such that
$\mathbb{E}\left[\sup_{t\in[0,T]}\|X^{\xi}_{t}-X_{t}^{\eta}\|_{\tau}^{k}\right]\leq c\text{\rm{e}}^{\varepsilon\left(\|\xi\|_{\tau}^{2\alpha}+\|\eta\|_{\tau}^{2\alpha}\right)}\|\xi-\eta\|_{\tau}^{k},\ \ \xi,\eta\in\mathscr{C}_{\tau}.$ (2.4)
2.2 Path-Distribution Dependent SDEs with Infinite Memory
To ensure the well-posedness of solutions to (1.2) with singular drift for distributions in $\mathscr{P}_{k}(\mathscr{C}_{\tau})$, we decompose $b$ as
$b(t,\xi,\mu)=b^{(0)}(t,\xi(0))+b^{(1)}(t,\xi,\mu),\ \ t\geq 0,\ \xi\in\mathscr{C}_{\tau},\ \mu\in\mathscr{P}_{k}(\mathscr{C}_{\tau}),$
and impose the following assumptions on $b^{(0)},b^{(1)}$ and $\sigma$.
  1. $(H_{1})$
    $b^{(0)}$ and $a:=\sigma\sigma^{*}$ satisfy $(A_{1})$ and $(A_{2})$.
  2. ($H_{2}$)
    $\sup_{t\geq 0}|b^{(1)}(t,{\bf 0},{\delta}_{\bf 0})|<\infty$, and there exists a function $H\in L^{1}_{loc}([0,\infty);(0,\infty))$ and constants $K>0$, $\alpha\in[0,1]$ such that
    $|b^{(1)}(t,\xi,\mu)-b^{(1)}(t,\eta,\nu)|^{2}\leq{K}\|\xi-\eta\|_{\tau}^{2}+H(t)\mathbb{W}_{k,var}(\mu,\nu)^{2},$ (2.5)
    $|b^{(1)}(t,\xi,\mu){-b^{(1)}(t,\xi^{0},\mu)}|\leq{K}\left\{1+\|\xi\|_{\tau}^{\alpha}+\|\mu\|_{k}\right\},$
    $\qquad\ \ t\geq 0,\ \xi,\eta\in\mathscr{C}_{\tau},\ \mu,\nu\in\mathscr{P}_{k}(\mathscr{C}_{\tau}),.$
Theorem 2.3.
[…]
$\mathscr{P}^{\alpha}_{k,\text{\rm{e}}}(\mathscr{C}_{\tau}):=\left\{\mu\in\mathscr{P}_{k}(\mathscr{C}_{\tau}):\mu\left(\text{\rm{e}}^{\varepsilon\|\cdot\|_{\tau}^{2\alpha}}\right)<\infty\text{ for some }\varepsilon\in(0,1)\right\}.$ (2.6)
[…]
$\mathbb{E}\left[\sup_{t\in[0,T]}\|X_{t}\|_{\tau}^{n}\bigg{|}X_{0}\right]\leq c\left(1+\|X_{0}\|_{\tau}^{n}\right).$ (2.7)
3 The Monotone Case: Well-Posedness and Asymptotic log-Harnack inequality
In this section, we consider (1.2) with monotone coefficients and establish the well-posedness and asymptotic log-Harnack inequality.
3.1 Path Dependent SDEs with Infinite Memory
Note that when the SDE
$\text{\rm{d}}X(t)=b(t,X_{t})\mathrm{d}t+\sigma(t,X_{t})\mathrm{d}W(t),\ \ t>0,\ \ X_{0}=\xi\in\mathscr{C}_{\tau}.$ (3.1)
is time-homogenous and monotone, the asymptotic log-Harnack inequality has been derived in [5, Theorem 3.1] for the Markov semigroup
$P_{t}f(\xi):=\mathbb{E}\big{[}f(X_{t}^{\xi})\big{]},\quad\ \ t\geq 0,\ \xi\in\mathscr{C}_{\tau},\ f\in\mathscr{B}_{b}(\mathscr{C}_{\tau}).$
We call $\mu$ an invariant probability measure of $P_{t}$, if it is a probability measure on $\mathscr{C}_{\tau}$ such that
$\int_{\mathscr{C}_{\tau}}P_{t}f\,\text{\rm{d}}\mu=\int_{\mathscr{C}_{\tau}}f\,\text{\rm{d}}\mu,\ \ f\in\mathscr{B}_{b}(\mathscr{C}_{\tau}).$
An important application of this inequality is the gradient estimate. For any function $f$ on $\mathscr{C}_{\tau}$, let
$|\nabla f(\xi)|:=\limsup_{\eta\rightarrow\xi}\frac{|f(\xi)-f(\eta)|}{\|\xi-\eta\|_{\tau}}$
be the Lipschitz constant of a function $f$ at point $\xi\in\mathscr{C}_{\tau}$, which coincides with the norm of the gradient $\nabla f(\xi)$ if $f$ is Gâteaux differentiable at $\xi$. We denote $f\in C_{b,L}(\mathscr{C}_{\tau})$ if
$\|f\|_{\infty}+\|\nabla f\|_{\infty}<\infty.$
Since the proof of [5, Theorem 3.1 and Theorem A.1] apply also to the time-inhomogenous case which is crucial in the study of distribution dependent SDEs, we reformulate this result and its consequence for (3.1) with time dependent coefficients without proof.
  1. $(B)$
    $b(t,\cdot)\in C(\mathscr{C}_{\tau}{;\mathbb{R}^{d}})$ for each $t\geq 0$, $\sigma$ is invertible, and there exists a constant $K>0$ such that
    $\|\sigma(t,\xi)-\sigma(t,\eta)\|^{2}+\langle b(t,\xi)-b(t,\eta),\xi(0)-\eta(0)\rangle^{+}\leq K\|\xi-\eta\|_{\tau}^{2},$
    $\|\sigma(t,\xi)\|+\|\sigma(t,\xi)^{-1}\|+|b(t,{\bf 0})|\leq K,\ \ \ t\geq 0,\ \xi,\eta\in\mathscr{C}_{\tau}.$
Theorem 3.1.
[…]
$\mathbb{E}\big{[}\|X_{t}^{\xi}\|_{\tau}^{2}\big{]}\leq c\text{\rm{e}}^{ct}\big{(}1+\|\xi\|_{\tau}^{2}\big{)},$
$P_{t}\log f(\eta)\leq\log P_{t}f(\xi)+c\|\xi-\eta\|_{\tau}^{2}+c\text{\rm{e}}^{-\tau_{0}t}\|\nabla\log f\|_{\infty}\|\xi-\eta\|_{\tau}.$ (3.2)
The following result is a direct consequence of Theorem 3.1 and [5, Theorem 2.1] with
$E=\mathscr{C}_{\tau},\ \ \ \rho(\xi,\eta)=\|\xi-\eta\|_{\tau},\ \ \ \Gamma_{t}(\xi)=c\text{\rm{e}}^{-\tau_{0}t},\ \ \ \Lambda(\xi)=c.$ (3.3)
Corollary 3.2.
In the situation of Theorem 3.1, the following assertions hold.
  1. $(1)$
    For any $t\geq 0,\xi\in\mathscr{C}_{\tau}$ and $f\in C_{b,L}(\mathscr{C}_{\tau})$,
    $|\nabla P_{t}f(\xi)|\leq\sqrt{2c[P_{t}f^{2}-(P_{t}f)^{2}]}(\xi)+c\text{\rm{e}}^{-\tau_{0}t}\|\nabla f\|_{\infty}.$
  2. $(2)$
    When the coefficients do not depend on $t$, $P_{t}$ has at most one invariant probability measure, and if $\mu$ is its invariant probability measure, then
    $\limsup_{t\rightarrow\infty}P_{t}f(\xi)\leq\log\left(\frac{\mu(\text{\rm{e}}^{f})}{\int_{\mathscr{C}_{\tau}}\text{\rm{e}}^{-c\|\xi-\eta\|_{\tau}^{2}}\mu(\text{\rm{d}}\eta)}\right),\ \ \xi\in\mathscr{C}_{\tau},\ f\in\mathscr{B}_{b}^{+}(\mathscr{C}_{\tau})\cap C_{b,L}(\mathscr{C}_{\tau}).$
    Consequently, for any closed set $A\subset\mathscr{C}_{\tau}$ with $\mu(A)=0$,
    $\lim_{t\rightarrow\infty}P_{t}{\bf 1}_{A}(\xi)=0.\ \ \xi\in\mathscr{C}_{\tau}.$
  3. $(3)$
    Let $\xi\in\mathscr{C}_{\tau}$ and $A\subset\mathscr{C}_{\tau}$ be a measurable set such that
    $\delta(\xi,A):=\liminf_{t\rightarrow\infty}P_{t}(\xi,A)>0.$
    Then for any $\varepsilon>0$ and $A_{\varepsilon}:=\{\eta\in\mathscr{C}_{\tau}:\|\eta-\xi\|_{\tau}<\varepsilon,\xi\in A\}$,
    $\liminf_{t\rightarrow\infty}P_{t}(\eta,A_{\varepsilon})>0,\ \ \eta\in\mathscr{C}_{\tau}.$
    More precisely, for any $\varepsilon_{0}\in(0,\delta(\xi,A))$, there exists a constant $t_{0}>0$ such that for any $\varepsilon>0$,
    $\inf\bigg{\{}P_{t}(\eta,A_{\varepsilon}):\ t>t_{0}\lor\Big{(}\frac{1}{\tau_{0}}\log\frac{\Lambda(\eta)\|\xi-\eta\|_{\tau}}{\varepsilon\varepsilon_{0}}\Big{)}\bigg{\}}>0,\ \ \eta\in\mathscr{C}_{\tau}.$
3.2 Path-Distribution Dependent SDEs with Infinite Memory
we assume that the following monotone assumption holds, as in [5] in the distribution-free case.
  1. $(H^{\prime})$
    $\sup_{t\geq 0}|b(t,{\bf 0},\delta_{\bf 0})|<\infty$, and $b(t,\cdot,\mu)\in C(\mathscr{C}_{\tau};\mathbb{R}^{d})$ holds for any $t\geq 0$, $\mu\in\mathscr{P}_{2}(\mathscr{C}_{\tau})$. Moreover, there exist constants $K_{1},K_{2}>0$ such that
    $\langle b(t,\xi,\mu)-b(t,\eta,\nu),\xi(0)-\eta(0)\rangle^{+}+\|\sigma(t,\xi)-\sigma(t,\eta)\|^{2}$
    $\leq K_{1}\|\xi-\eta\|_{\tau}^{2}+K_{2}\mathbb{W}_{2}(\mu,\nu)^{2},\ \ t\geq 0,\ \xi,\eta\in\mathscr{C}_{\tau},\ \mu,\nu\in\mathscr{P}_{2}(\mathscr{C}_{\tau}).$
Theorem 3.3.
Assume $(H^{\prime})$. Then the following assertions hold.
  1. $(1)$
    The SDE (1.2) is well-posed for distributions in $\mathscr{P}_{2}(\mathscr{C}_{\tau})$, and there exists a constant $c>0$ such that
    $\mathbb{W}_{2}(P_{t}^{*}\mu_{0},P_{t}^{*}\nu_{0})\leq c\text{\rm{e}}^{ct}\mathbb{W}_{2}(\mu_{0},\nu_{0}),\ \ t\geq 0,\ \mu_{0},\nu_{0}\in\mathscr{P}_{2}(\mathscr{C}_{\tau}).$ (3.4)
  2. $(2)$
    If $\sigma$ is invertible with $\|\sigma\|_{\infty}+\|\sigma^{-1}\|_{\infty}<\infty$, then for any $\tau_{0}\in(0,\tau)$, there exists a constant $c\geq 1$ such that
    $P_{t}\log f(\nu)\leq$ $\log P_{t}f(\mu)+c_{0}(t)\mathbb{W}_{2}(\mu,\nu)^{2}+c\text{\rm{e}}^{-\tau_{0}t}\|\nabla\log f\|_{\infty}\mathbb{W}_{2}(\mu,\nu)$ (3.5)
    holds for $t\geq 0$, $\mu,\nu\in\mathscr{P}_{2}(\mathscr{C}_{\tau})$ and $f\in\mathscr{B}^{+}_{b}(\mathscr{C}_{\tau})$ with $\|\nabla\log f\|_{\infty}<\infty$, where $c_{0}(t):=c\left(K_{2}\text{\rm{e}}^{ct}+1\right)$. Consequently, for any $t>0$ and $f\in C_{b,L}(\mathscr{C}_{\tau})$,
    $|P_{t}f(\mu)-P_{t}f(\nu)|\leq\mathbb{W}_{2}(\mu,\nu)\big{[}c\text{\rm{e}}^{-\tau_{0}t}\|\nabla f\|_{\infty}+{2\sqrt{C(K_{2},t)}}\|f\|_{\infty}\big{]}.$ (3.6)
Remark 3.4.
Note that (3.5) implies (3.1) for $\mu=\delta_{\xi}$, $\nu=\delta_{\eta}$ and $K_{2}=0$.
4 Proofs of Theorem 2.1 Theorem 2.2
To prove these two theorems , we present some lemmas below.
Lemma 4.1.
Assume $(A_{1})$, $(A_{2})$, $(A_{3}^{\prime})$ and $\|b^{(1)}\|_{\infty}<\infty$. Then for any initial value $\xi\in\mathscr{C}_{\tau}$, (2.1) has a unique non-explosive strong solution satisfying (2.3), and for any $T\in(0,\infty)$ there exists a constant $c>0$ such that
${\mathbb{E}\left[\sup_{t\in[0,T]}\big{|}X^{\xi}(t)-X^{\eta}(t)\big{|}\right]\leq c\|\xi-\eta\|_{\tau}},\ \ \ \xi,\eta\in\mathscr{C}_{\tau}.$ (4.1)
Proof.
The desired estimate follows from the proof of [18, Lemma 4.1] with $\|\cdot\|_{\tau}$ in place of uniform norm $\|\cdot\|_{\mathscr{C}}$ on the path space with finite time interval. ∎
Next, consider the local Hardy-Littlewood maximal function for a nonnegative function $f$ on $\mathbb{R}^{d}$:
$\mathscr{M}f(x):=\sup_{r\in(0,1)}\frac{1}{|B(0,r)|}\int_{B(0,r)}f(x+y)\,\mathrm{d}y,\ \ x\in\mathbb{R}^{d},$ (4.2)
where $|B(0,r)|$ is the volume of $B(0,r):=\{y:|y|<r\}$. The following result is taken from [15, Lemma 2.1].
Lemma 4.2.
  1. $(1)$
    There exists a constant $c>0$ such that for any $f\in C_{b}(\mathbb{R}^{d})$ with $|\nabla f|\in L^{1}_{loc}(\mathbb{R}^{d})$,
    $|f(x)-f(y)|\leq c|x-y|(\mathscr{M}|\nabla f|(x)+\mathscr{M}|\nabla f|(y)+\|f\|_{\infty}),\ \ x,y\in\mathbb{R}^{d}.$ (4.3)
  2. $(2)$
    For any $T,p,q\in(1,\infty),$ there exists a constant $c>0$ such that
    $\|\mathscr{M}f\|_{\tilde{L}_{q}^{p}(T)}\leq c\|f\|_{\tilde{L}_{q}^{p}(T)},\ \ f\in\tilde{L}_{q}^{p}(T).$ (4.4)
The lemma below examines the exponential integrability of functionals for segment process.
Lemma 4.3.
Assume $(A_{1})$, $(A_{2})$ and $(A_{3}^{\prime})$. For any $T\in(1,\infty)$, there exist constants $\beta_{T},k_{T}>0$ such that for any solution $(X_{t})_{t\in[0,T]}$ to (2.1),
$\mathbb{E}\left[\exp\left(\beta\int_{s}^{t}\|X_{u}\|_{\tau}^{2\alpha}\,\mathrm{d}u\right)\bigg{|}\mathscr{F}_{s}\right]\leq\text{\rm{e}}^{k_{T}\beta\left(1+\|X_{s}\|_{\tau}^{2\alpha}\right)}$ (4.5)
holds for any $0\leq\beta\leq\beta_{T}$ and $0\leq s\leq t\leq T$.
Proof.
Obviously, (4.5) holds for $\alpha=0$. By shifting the starting time from 00 to $s$, we may and do assume that $s=0$. So, by Jensen’s inequality, it suffices to find constants $\beta_{T},k_{T}>0$ such that
$\mathbb{E}\left[\exp\left(\beta_{T}\int_{0}^{T}\|X_{u}\|_{\tau}^{2\alpha}\,\mathrm{d}u\right)\bigg{|}\mathscr{F}_{0}\right]\leq\text{\rm{e}}^{k_{T}{\beta_{T}}\left(1+\|X_{0}\|_{\tau}^{2\alpha}\right)}.$ (4.6)
By [14, Theorem 3.2] and [17, Theorem 2.1], there exists a constant $\lambda_{0}>0$ such that for any $\lambda\geq\lambda_{0}$, the PDE for $u(t,\cdot):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$
$(\partial_{t}+L^{0}(t))u(t,\cdot)=\lambda u(t,\cdot)-b^{(0)}(t,\cdot),\ \ t\in[0,T],\ \ u_{T}=0$ (4.7)
has a unique solution in $\tilde{H}_{q_{0}}^{2,p_{0}}(T)$, where $L^{0}(t):=\nabla_{b^{(0)}(t,\cdot)}+\frac{1}{2}\text{tr}(\{\sigma\sigma^{*}\}(t,\cdot)\nabla^{2})$, and there exist constants $c,\theta>0$ such that
$\lambda^{\theta}(\|u\|_{\infty}+\|\nabla u\|_{\infty})+\|\partial_{t}u\|_{\tilde{L}_{q_{0}}^{p_{0}}(T)}+\|\nabla^{2}u\|_{\tilde{L}_{q_{0}}^{p_{0}}(T)}\leq c,\ \ \lambda\geq\lambda_{0}.$ (4.8)
We may take $\lambda\geq\lambda_{0}$ such that
$\|u\|_{\infty}+\|\nabla u\|_{\infty}\leq\frac{1}{2},$ (4.9)
so that $\Theta(t,x):=x+u(t,x)$ satisfies
$\frac{1}{2}|x-y|^{2}\leq|\Theta(t,x)-\Theta(t,y)|^{2}\leq 2|x-y|^{2},\ \ t\in[0,T],\ x,y\in\mathbb{R}^{d}.$ (4.10)
For any $t\in[0,T]$, let $Y(t)=\Theta(t,X(t))$. By Itô’s formula in [10, Theorem 1.2.3(3)] and (4.7), we obtain
$\text{\rm{d}}Y(t)=\big{\{}\lambda u(t,X(t))+\nabla\Theta(t,X(t))b^{(1)}(t,X_{t})\big{\}}\text{\rm{d}}t+\big{(}\{\nabla\Theta\}\sigma\big{)}(t,X(t))\text{\rm{d}}W(t).$
Combining this with $(A_{1})$, $(A_{2})$, the second inequality in $(A_{3}^{\prime})$, (4.9) and Itô’s formula, we find a constant $c_{0}>0$ such that
$\text{\rm{d}}\{|Y(t)|^{2}+1\}^{\alpha}\leq$ $c_{0}\{|Y(t)|^{2}+1\}^{\alpha-\frac{1}{2}}\{1+|Y(t)|+\|Y_{t}\|^{\alpha}_{\tau}\}\text{\rm{d}}t+\text{\rm{d}}M(t),\ \ t\in[0,T],$ (4.11)
where
$\text{\rm{d}}M(t)=2\alpha\{|Y(t)|^{2}+1\}^{(\alpha-1)}\langle Y(t),\big{(}\{\nabla\Theta\}\sigma\big{)}(t,X(t))\text{\rm{d}}W(t)\rangle.$
Note that
$\begin{split}&\text{\rm{e}}^{p\tau t}\|X_{t}\|_{\tau}^{p}:=\sup_{s\in(-\infty,0]}\left(\text{\rm{e}}^{p\tau(t+s)}|X(t+s)|^{p}\right)\\ &\leq\sup_{s\in(-\infty,0]}\left(\text{\rm{e}}^{p\tau s}|X(s)|^{p}\right)+\sup_{s\in[0,t]}\left(\text{\rm{e}}^{p\tau s}|X(s)|^{p}\right)\\ &=\|X_{0}\|_{\tau}^{p}+\sup_{s\in[0,t]}\left(\text{\rm{e}}^{p\tau s}|X(s)|^{p}\right),\ \ t\geq 0{,\,p\geq 0.}\end{split}$ (4.12)
Combining this with (4.11), Young’s inequality, Itô’s formula, and $(s+1)^{r}\leq 1$ for $s\geq 0$ and $r\leq 0$, we find a constant $c_{1}>0$ such that
$\begin{split}\text{\rm{d}}\big{\{}\text{\rm{e}}^{2\alpha\tau t}\{|Y(t)|^{2}+1\}^{\alpha}\big{\}}&\leq\text{\rm{e}}^{2\alpha\tau t}\text{\rm{d}}M(t)+c_{1}\left(1+\|Y_{0}\|_{\tau}^{2\alpha}\right)\text{\rm{d}}t\\ &+\sup_{s\in[0,t]}\text{\rm{e}}^{2\alpha\tau s}\left(|Y(s)|^{2}+1\right)^{\alpha}\text{\rm{d}}t,\ \ t\in[0,T].\end{split}$ (4.13)
So, letting
$l(t)=\text{\rm{e}}^{2\alpha\tau t}\{|Y(t)|^{2}+1\}^{\alpha}+c_{1}\left(1+\|Y_{0}\|_{\tau}^{2\alpha}\right),$
$\bar{l}_{t}=\sup_{s\in[0,t]}l(s)\ \ t\in[0,T],$
we find a constant $c>0$ and a martingale $\tilde{M}(t)$ such that
$\text{\rm{d}}l(t)\leq c\,\bar{l}_{t}\text{\rm{d}}t+\text{\rm{d}}\tilde{M}(t).$
Then by [6, Lemma 3.1], we find a constant $c_{2}>0$ such that
$\mathbb{E}\left[\exp\bigg{(}\frac{\varepsilon}{T\text{\rm{e}}^{1+cT}}\int_{0}^{t}\bar{l}_{s}\,\text{\rm{d}}s\bigg{)}\bigg{|}\mathscr{F}_{0}\right]\leq\text{\rm{e}}^{\varepsilon l(s)+1}\left(\mathbb{E}\left[\exp\left(2\varepsilon^{2}\langle\tilde{M}\rangle(t)\right)\big{|}\mathscr{F}_{0}\right]\right)^{\frac{1}{2}}$
$\leq\text{\rm{e}}^{\varepsilon l(0)+1}\left(\mathbb{E}\left[\exp\left(c_{2}\varepsilon^{2}\int_{0}^{t}l(s)\,\text{\rm{d}}s\right)\bigg{|}\mathscr{F}_{0}\right]\right)^{\frac{1}{2}},\ \ \varepsilon>0,\ t\in[0,T].$
Taking $\varepsilon=\frac{1}{c_{2}T\text{\rm{e}}^{1+cT}},$ we derive
$\mathbb{E}\left[\exp\bigg{(}\frac{\varepsilon}{T\text{\rm{e}}^{1+cT}}\int_{0}^{t}\bar{l}_{s}\,\text{\rm{d}}s\bigg{)}\bigg{|}\mathscr{F}_{0}\right]\leq\text{\rm{e}}^{2(\varepsilon l(0)+1)},\ \ t\in[0,T].$
By combining this with (4.9) and (4.12), we derive (4.6) for $\beta_{T}=\frac{\varepsilon}{T\text{\rm{e}}^{1+cT}}$ and some constant $k_{T}>0$. ∎
In the next result, we present Krylov-Khasminskii estimates for SDEs with memory.
Lemma 4.4.
Assume $(A_{1})$, $(A_{2})$ and $(A_{3}^{\prime})$. For any constants $T,p,q>1$ such that $(2p,2q)\in\mathscr{K}$ and $\varepsilon>0$, there exist constants $c=c(T,p,q,\varepsilon)>0$ and $\theta=\theta(T,p,q)>0$ such that the solution to (2.1) satisfies
$\mathbb{E}\left[\int_{s}^{t}|f_{u}(X(u))|\,\text{\rm{d}}u\bigg{|}\mathscr{F}_{s}\right]\leq\left(c+\varepsilon\|X_{s}\|_{\tau}^{\alpha}\right)\|f\|_{\tilde{L}_{q}^{p}(s,t)},$ (4.14)
$\mathbb{E}\left[\exp\left(\int_{s}^{t}|f_{u}(X(u))|\,\text{\rm{d}}u\right)\bigg{|}\mathscr{F}_{s}\right]\leq\exp\left[\varepsilon{\|X_{s}\|}_{\tau}^{2\alpha}+c\left(1+\|f\|^{\theta}_{\tilde{L}^{p}_{q}(s,t)}\right)\right]$ (4.15)
for any $0\leq s\leq t\leq T$ and $f\in\tilde{L}^{p}_{q}(T)$.
Proof.
Without loss of generality, we may simply prove these estimates for $s=0$.
For any $x\in\mathbb{R}^{d}$, let $\phi^{x}\in\mathscr{C}_{\tau}$ be defined as
$\phi^{x}(r):=x,\ \ r\leq 0.$
For any $t\in[0,T],\ x\in\mathbb{R}^{d}$ and $\xi\in\mathscr{C}_{\tau}$, let
$\bar{b}(t,x):=b^{(0)}(t,x)+b^{(1)}(t,\phi^{x}),\ \ \ \hat{b}(t,\xi):=b^{(1)}(t,\xi)-b^{(1)}(t,\xi^{0}).$
By (4.9), (4.12), (4.13) and the Burkholder-Davis-Gundy (BDG) inequality, we find a constant $k_{1}>0$ such that
$\mathbb{E}\left[\sup_{t\in[0,T]}\|X_{t}\|_{\tau}^{2\alpha}\bigg{|}\mathscr{F}_{0}\right]\leq k_{1}\left(1+\|X_{0}\|_{\tau}^{2\alpha}\right).$ (4.16)
For any $n\geq 1$, let
$\tau_{n}:=\inf\{t\in[0,T]:\|X_{t}\|_{\tau}\geq n\},\ \ \inf\emptyset:=T.$ (4.17)
By (4.16), $(A_{1})$, $(A_{2})$, the second inequality in $(A_{3}^{\prime})$, and applying Krylov’s estimate in [17, Theorem 3.1] for $\bar{b}(t,x)$ in place of $b(t,x)$, see also [10, Theorem 1.2.3(2)], we find constants $k_{2}$ depending on $\varepsilon$ and $k_{3}>0$ independent of $\varepsilon$ such that
$\mathbb{E}\left[\int_{0}^{t\wedge\tau_{n}}|f_{s}|\left(X(s)\right)\,\text{\rm{d}}s\bigg{|}\mathscr{F}_{0}\right]$
$\leq$ $\left\{k_{2}+\varepsilon\left[\mathbb{E}\left(\int_{0}^{t\wedge\tau_{n}}|\hat{b}(s,X_{s})|^{2}\,\text{\rm{d}}s\bigg{|}\mathscr{F}_{0}\right)\right]^{\frac{1}{2}}\right\}\|f\|_{\tilde{L}_{q}^{p}(0,t)}$
$\leq$ $k_{3}\left(k_{2}+\varepsilon\|X_{0}\|_{\tau}^{\alpha}\right)\|f\|_{\tilde{L}_{q}^{p}(0,t)},\ \ t\in[0,T].$
Letting $n\rightarrow\infty$, we derive (4.14) since $\varepsilon>0$ is arbitrary.
To prove (4.15), we shall use Girsanov’s transform. Let
$\gamma_{t}=\{\sigma^{*}(\sigma\sigma^{*})^{-1}\}(t,X(t))\hat{b}(t,X_{t}),\ \ t\in[0,T].$ (4.18)
Then for any $n\geq 1$ and $\tau_{n}$ in (4.17),
$R_{n}(t):=\exp\bigg{[}-\int_{0}^{t\land\tau_{n}}\langle\gamma_{r},\text{\rm{d}}W(r)\rangle-\frac{1}{2}\int_{0}^{t\land\tau_{n}}|\gamma_{r}|^{2}\,\text{\rm{d}}r\bigg{]},\ \ t\in[0,T]$
is a martingale. By Girsanov’s theorem,
${W}_{n}(t):=W(t)+\int_{0}^{t\wedge\tau_{n}}\gamma_{s}\,\text{\rm{d}}s,\ \ t\in[0,T]$
is a Brownian motion under the probability measure $\text{\rm{d}}\mathbb{Q}_{n}:=R_{n}(T)\text{\rm{d}}\mathbb{P}$, and $X$ solves the equation
$X(t)=X(0)+\int_{0}^{t}\bar{b}(s,X(s))\,\text{\rm{d}}s+\int_{0}^{t}\sigma(s,X(s))\,\text{\rm{d}}W_{n}(s),\ \ t\in[0,\tau_{n}].$ (4.19)
Since the Krylov estimate in [10, Theorem 1.2.4] holds for (4.19), letting $f^{n}_{r}(x)=f_{r}(x){\bf 1}_{\{|x|<n\}}$, by the argument in the proof of [18, Lemma 4.4], we find a constant $k>0$ independent of $s$ such that for any $t\in[0,T]$,
$\mathbb{E}_{\mathbb{Q}_{n}}\left[\exp\left(\lambda\int_{0}^{t}\left|f^{n}_{r}\left(X(r\wedge\tau_{n})\right)\right|\,\text{\rm{d}}r\right)\bigg{|}\mathscr{F}_{0}\right]\leq\exp\Big{[}k+k\|f\|_{\tilde{L}_{q}^{p}(0,t)}^{k}\Big{]}.$ (4.20)
On the other hand, for constants $K$ in $(A_{1})$ $(A_{2})$, $(A_{3}^{\prime})$, and $\beta_{T}$ in Lemma 4.3, let
$p_{0}:=\frac{3+\sqrt{1+8\beta_{T}K^{-3}}}{4}.$
We have
$k(p):=(p-1)(2p-1)K^{3}\in(0,\beta_{T}],\ \ \ p\in(1,p_{0}].$
Thus, taking $\tilde{p}\in[1,p_{0}]$ with $\tilde{p}-1$ small enough such that ${k_{T}}k(\tilde{p})\leq 2\varepsilon$, by $(A_{1})$ $(A_{2})$, $(A_{3}^{\prime})$ and (4.5), we find a constant $\beta=\beta(p,q,\varepsilon)>0$ such that
$\mathbb{E}\big{[}|R_{n}(T)|^{-(\tilde{p}-1)}\big{|}\mathscr{F}_{0}\big{]}$ (4.21)
$=$ $\mathbb{E}\left[\exp\bigg{(}(\tilde{p}-1)\int_{0}^{T\wedge\tau_{n}}\langle\gamma_{s},\text{\rm{d}}W(s)\rangle+\frac{\tilde{p}-1}{2}\int_{0}^{T\wedge\tau_{n}}|\gamma_{s}|^{2}\,\text{\rm{d}}s\bigg{)}\bigg{|}\mathscr{F}_{0}\right]$
$\leq$ $\left(\mathbb{E}\left[\exp\bigg{(}(\tilde{p}-1)(2\tilde{p}-1)\int_{0}^{T\wedge\tau_{n}}|\gamma_{s}|^{2}\,\text{\rm{d}}s\bigg{)}\bigg{|}\mathscr{F}_{0}\right]\right)^{\frac{1}{2}}$
×$\quad\times\left(\mathbb{E}\left[\exp\bigg{(}2(\tilde{p}-1)\int_{0}^{T\wedge\tau_{n}}\langle\gamma_{s},\text{\rm{d}}W(s)\rangle-2(\tilde{p}-1)^{2}\int_{0}^{T\wedge\tau_{n}}|\gamma_{s}|^{2}\,\text{\rm{d}}s\bigg{)}\bigg{|}\mathscr{F}_{0}\right]\right)^{\frac{1}{2}}$
$\leq$ $\exp\left[\beta+\varepsilon\|X_{0}\|^{2\alpha}_{\tau}\right]\times 1=\exp\left[\beta+\varepsilon\|X_{0}\|^{2\alpha}_{\tau}\right].$
Combining this with (4.20) and applying Hölder’s inequality, we find constants $c=c(p,q,\varepsilon)>0$ and $\theta=\theta(p,q)>0$ such that
$\mathbb{E}\left[\exp\left(\int_{0}^{t}\left|f^{n}_{s}(X(s\wedge\tau_{n}))\right|\,\text{\rm{d}}s\right)\bigg{|}\mathscr{F}_{0}\right]$
$=$ $\mathbb{E}_{\mathbb{Q}_{n}}\left[R_{n}(T)^{-1}\exp\left(\int_{0}^{t}|f^{n}_{s}(X(s\wedge\tau_{n}))|\,\text{\rm{d}}s\right)\bigg{|}\mathscr{F}_{0}\right]$
$\leq$ $\left(\mathbb{E}\big{[}R_{n}(T)^{-(\tilde{p}-1)}\big{|}\mathscr{F}_{0}\big{]}\right)^{\frac{1}{\tilde{p}}}\left(\mathbb{E}_{\mathbb{Q}_{n}}\left[\exp\left(\frac{\tilde{p}}{\tilde{p}-1}\int_{0}^{t}|f^{n}_{r}(X(s\wedge\tau_{n}))|\,\text{\rm{d}}s\right)\bigg{|}\mathscr{F}_{0}\right]\right)^{\frac{\tilde{p}-1}{\tilde{p}}}$
$\leq$ $\exp\left[\varepsilon\|X_{0}\|_{\tau}^{2\alpha}+c\left(1+\|f\|^{\theta}_{\tilde{L}_{p}^{q}(0,t)}\right)\right].$
This with $n\rightarrow\infty$ implies (4.15). ∎
By Lemma 4.3 and (4.15), we have the following corollary which is a key in the proofs of Theorem 2.2 and 2.3.
Corollary 4.5.
Assume $(A_{1})$, $(A_{2})$ and $(A_{3}^{\prime})$.
  1. $(1)$
[…]
    $\mathbb{E}\left[\exp\left(\int_{s}^{t}|f_{u}(X(u))|\|X_{u}\|_{\tau}^{\alpha}\,\text{\rm{d}}u\right)\bigg{|}\mathscr{F}_{s}\right]\leq\exp\left[\varepsilon{\|X_{s}\|}_{\tau}^{2\alpha}+k\left(1+\|f\|^{\theta}_{\tilde{L}^{p}_{q}(s,t)}\right)\right],$ (4.22)
[…]
  2. $(2)$
[…]
    $\sup_{t\in[0,T]}\mathbb{E}\Big{[}\text{\rm{e}}^{\beta\|X_{t}\|_{\tau}^{2\alpha}}\Big{|}\mathscr{F}_{0}\Big{]}\leq\text{\rm{e}}^{k_{T}\beta(1+\|X_{0}\|_{\tau}^{2\alpha})},\ \ \beta\in[0,\beta_{T}].$ (4.23)
Proof.
(a) Let $\beta_{T},k_{T}>0$ be in Lemma 4.3, and let $\varepsilon,c>0$ be in Lemma 4.4. We choose $\beta\in(0,\beta_{T}]$ such that $k_{T}\beta\leq 2\varepsilon$, so that by Lemma 4.3, (4.15) and Young’s and Hölder’s inequalities, we find constants $\theta=\theta(T,p,q),k=k(T,p,q,\varepsilon)>0$ such that
$\mathbb{E}\left[\exp\left(\int_{s}^{t}|f_{u}(X(u))|\cdot\|X_{u}\|_{\tau}^{\alpha}\,\text{\rm{d}}u\right)\bigg{|}\mathscr{F}_{s}\right]$
$\leq$ $\mathbb{E}\left[\exp\left(\int_{s}^{t}\left({\frac{1}{2\beta}}|f_{u}(X(u))|^{2}+\frac{\beta}{2}\cdot\|X_{u}\|_{\tau}^{2\alpha}\right)\,\text{\rm{d}}u\right)\bigg{|}\mathscr{F}_{s}\right]$
$\leq$ $\left(\mathbb{E}\left[\exp\left({\frac{1}{\beta}}\int_{s}^{t}|f_{u}(X(u))|^{2}\,\text{\rm{d}}u\right)\bigg{|}\mathscr{F}_{s}\right]\right)^{\frac{1}{2}}\left(\mathbb{E}\left[\exp\left(\beta\int_{s}^{t}\|X_{u}\|_{\tau}^{2\alpha}\,\text{\rm{d}}u\right)\bigg{|}\mathscr{F}_{s}\right]\right)^{\frac{1}{2}}$
$\leq$ $\exp\left[\varepsilon{\|X_{s}\|}_{\tau}^{2\alpha}+k\left(1+{\|f^{2}\|^{\theta/2}_{\tilde{L}^{p/2}_{q/2}(s,t)}}\right)\right]=\exp\left[\varepsilon{\|X_{s}\|}_{\tau}^{2\alpha}+k\left(1+\|f\|^{\theta}_{\tilde{L}^{p}_{q}(s,t)}\right)\right].$
(b) In the case where $b^{(1)}=0$, (4.23) can be obtained by repeating the proof of [4, Lemma 2.4] and [3, Corollary 2.5], so that by Girsanov’s theorem, for the segment solution to (4.19) and $\mathbb{Q}_{n}$ in the proof of Lemma 4.4, there exists constants $c_{0},c_{1}>0$ such that
$\mathbb{E}_{\mathbb{Q}_{n}}\left[\text{\rm{e}}^{c_{0}\|X_{t\land\tau_{n}}\|_{\tau}^{2\alpha}}\big{|}\mathscr{F}_{0}\right]\leq\text{\rm{e}}^{c_{1}+c_{1}\|X_{0}\|_{\tau}^{2\alpha}},\ \ \ t\in[0,T],\ n\geq 1.$ (4.24)
Letting $\tilde{p}>1$ be in (4.21) and taking
$\beta_{T}:=\frac{\tilde{p}c_{0}}{\tilde{p}-1},$
by combining (4.21) with (4.24) and using Hölder’s inequality, we find a constant $c>0$ such that
$\mathbb{E}\big{[}\text{\rm{e}}^{\beta_{T}\|X_{t\land\tau_{n}}\|_{\tau}^{2\alpha}}\big{|}\mathscr{F}_{0}\big{]}\leq\mathbb{E}_{\mathbb{Q}_{n}}\left[R_{n}(T)^{-1}\text{\rm{e}}^{\beta_{T}\|X_{t{\land\tau_{n}}}\|_{\tau}^{2\alpha}}\big{|}\mathscr{F}_{0}\right]$
$\leq$ $\left(\mathbb{E}\big{[}R_{n}(T)^{-(\tilde{p}-1)}\bigg{|}\mathscr{F}_{0}\big{]}\right)^{\frac{1}{\tilde{p}}}\left(\mathbb{E}_{\mathbb{Q}_{n}}\left[\exp\left(\frac{\tilde{p}\beta_{T}}{\tilde{p}-1}\|X_{t{\land\tau_{n}}}\|_{\tau}^{2\alpha}\right)\bigg{|}\mathscr{F}_{0}\right]\right)^{\frac{\tilde{p}-1}{\tilde{p}}}$
$\leq$ $\exp\left(c+c\|{X_{0}}\|_{\tau}^{2\alpha}\right),\ \ t\in[0,T],\ n\geq 1.$
By letting $n\rightarrow\infty$ we derive
$\mathbb{E}\big{[}\text{\rm{e}}^{\beta_{T}\|X_{t}\|_{\tau}^{2\alpha}}\big{|}\mathscr{F}_{0}\big{]}\leq\exp\left(c+c\|{X_{0}}\|_{\tau}^{2\alpha}\right),\ \ t\in[0,T],$
which implies (4.23) for ${k}_{T}:=\frac{c}{\beta_{T}}$ by Jensen’s inequality. ∎
Proof of Theorem 2.1.
Let $X_{0}$ be an $\mathscr{F}_{0}$-measurable random variable in $\mathscr{C}_{\tau}$. Take $\psi\in C_{b}^{\infty}([0,\infty))$ such that $0\leq\psi\leq 1$, $\psi(u)=1$ for $u\in[0,1]$, and $\psi(u)=0$ for $u\in[2,\infty)$. Define
$b^{(1)}_{n}(t,\xi)=b^{(1)}(t,\xi)\psi\left(\frac{\|\xi\|_{\tau}}{n}\right),\ \ t\geq 0,\ \xi\in\mathscr{C}_{\tau},\,n\geq 1.$ (4.25)
By (2.2), for any $n\geq 1$, $b^{(1)}_{n}$ is bounded and satisfies $(A_{3}^{\prime})$. So, Lemma 4.1 ensures that the SDE
$\text{\rm{d}}X^{n}(t)=b^{(1)}_{n}(t,X^{n}_{t})\text{\rm{d}}t+b^{(0)}(t,X^{n}(t))\text{\rm{d}}t+\sigma(t,X^{n}(t))\text{\rm{d}}W(t),\ \ X_{0}^{n}=X_{0},\ t\in[0,T]$
has a unique strong solution. Let $\tau_{n}$ be in (4.17). It suffices to prove that for any $k\geq 2$ there exists a constant $c(T,k)>0$ such that
$\mathbb{E}\bigg{[}\sup_{t\in[0,T]}\|X_{t}^{n}\|_{\tau}^{k}\bigg{|}\mathscr{F}_{0}\bigg{]}\leq c(T,k)\big{(}1+\|X_{0}\|_{\tau}^{k}\big{)},\ \ n\geq 1.$ (4.26)
This implies $\mathbb{P}(\tau_{n}=T)\rightarrow 1$ as $n\rightarrow\infty$ and
$X(t):=1_{\{t=0\}}X(0)+\sum_{n=1}^{\infty}X_{n}(t)1_{(\tau_{n-1},\tau_{n}]}(t),\ \ t\in[0,T]$
is the unique solution to (2.1) and (2.3) holds, where $\tau_{0}:=0$. To prove (4.26), we use Zvonkin’s transform.
For $\lambda\geq\lambda_{0}$ and $u$ solving the PDE (4.7) such that (4.8) and (4.9) hold, let
$\Theta(t,x):=x+u(t,x),\ \ Y^{n}(t):=\Theta(t,X^{n}(t)),\ \ t\in[0,T].$
By Itô’s formula, we obtain
$\text{\rm{d}}Y^{n}(t)=$$\{\nabla\Theta(t,X^{n}(t))b^{(1)}_{n}(t,X_{t}^{n})+\lambda u(t,X^{n}(t))\}\text{\rm{d}}t$ (4.27)
+$+\{(\nabla\Theta\sigma)(t,X^{n}(t))\}\text{\rm{d}}W(t).$
Let
$Z(t)=\text{\rm{e}}^{k\tau t}|Y^{n}(t)|^{k}+1.$
By $(A_{1})$$(A_{3})$, (4.9) and Itô’s formula, we find a constant $c_{1}>0$ such that
$\text{\rm{d}}Z(t)\leq$ $c_{1}\left(1+\|Y^{n}_{0}\|_{\tau}^{k}\right)\text{\rm{d}}t+c_{1}\sup_{s\in[0,t]}Z(s)\text{\rm{d}}t+\text{\rm{d}}M(t),\ \ t\leq\tau_{n},$ (4.28)
where
$\text{\rm{d}}M(t)=k\text{\rm{e}}^{k\tau t}|Y^{n}(t)|^{k-2}\langle Y^{n}(t),\big{(}\{\nabla\Theta\}\sigma\big{)}(t,X^{n}(t))\text{\rm{d}}W(t)\rangle.$
By the BDG inequality and the Young inequality, we find constants $c_{2},c_{3}>0$ such that
$\mathbb{E}\bigg{[}\sup_{s\in[0,t\land\tau_{n}]}M(s)\bigg{|}\mathscr{F}_{0}\bigg{]}\leq c_{2}\mathbb{E}\bigg{[}\int_{0}^{t}Z(s\land\tau_{n})^{2}\,\text{\rm{d}}s\bigg{|}\mathscr{F}_{0}\bigg{]}^{\frac{1}{2}}$
$\leq$ $\frac{1}{2}\mathbb{E}\bigg{[}\sup_{s\in[0,t\land\tau_{n}]}Z(s)\bigg{|}\mathscr{F}_{0}\bigg{]}+c_{3}\int_{0}^{t}\mathbb{E}\big{[}Z(s\land\tau_{n})\big{|}\mathscr{F}_{0}\big{]}\,\text{\rm{d}}s,\ \ t\in[0,T].$
Combining this with (4.12), (4.28) and
$|X^{n}(t)-Y^{n}(t)|\leq\|u\|_{\infty}<\infty,$
we may apply Gronwall’s inequality to derive the desired (4.26) for some constant $c=c(T,k)>0.$
Proof of Theorem 2.2.
Let $\Theta(t,x):=x+u(t,x)$ for $u$ solving the PDE (4.7) such that (4.8) and (4.9) hold, and denote
$Y^{\xi}(t):=\Theta(t,X^{\xi}(t)),\ \ \ Y^{\eta}(t):=\Theta(t,X^{\eta}(t)),$
$\Phi(t):=\text{\rm{e}}^{\tau t}(X^{\xi}(t)-X^{\eta}(t)),\ \ \Xi(t):=\text{\rm{e}}^{2k\tau t}|Y^{\xi}(t)-Y^{\eta}(t)|^{2k},$
$g_{1}(t):=\mathscr{M}\|\nabla^{2}u\|(t,X^{\xi}(t))+\mathscr{M}\|\nabla^{2}u\|(t,X^{\eta}(t)),$
$g_{2}(t):=g_{1}(t)+\mathscr{M}\|\nabla\sigma\|(t,X^{\xi}(t))+\mathscr{M}\|\nabla\sigma\|(t,X^{\eta}(t)),$
$A(t):=\int_{0}^{t}\left(1+g_{2}(s)^{2}+g_{1}(s)\|X^{\xi}_{s}\|_{\tau}^{\alpha}\right)\text{\rm{d}}s,\ \ \ t\in[0,T].$
By Lemma 4.2, Lemma 4.4, Corollary 4.5 and Holder’s inequality, for any $\beta>0$ and $\varepsilon\in(0,1)$ we find a constants $c(\beta,\varepsilon)>0$ such that
$\mathbb{E}\left[\text{\rm{e}}^{\beta{A}(T)}\right]\leq\text{\rm{e}}^{c(\beta,\varepsilon)+\varepsilon(\|\xi\|_{\tau}^{2\alpha}+\|\eta\|_{\tau}^{2\alpha})}.$ (4.29)
By $(A_{1})$, $(A_{2})$, $(A_{3}^{\prime})$, (4.9) and Lemma 4.2, we find a constant $k_{0}>0$ such that
$\left|[\nabla\Theta(t,X^{\xi}(t))]b^{(1)}(t,X^{\xi}_{t})-[\nabla\Theta(t,X^{\eta}(t))]b^{(1)}(t,X^{\eta}_{t})\right|$
$\leq k_{0}\|X^{\xi}_{t}-X_{t}^{\eta}\|_{\tau}+k_{0}\left(1+g_{1}(t)\right)\|X_{t}^{\xi}\|_{\tau}^{\alpha}|X^{\xi}(t)-X^{\eta}(t)|,$
$\left\|\left([\nabla\Theta]\sigma\right)(t,X^{\xi}(t))-\left([\nabla\Theta]\sigma\right)(t,X^{\eta}(t))\right\|^{2}_{HS}\leq k_{0}\left(1+g_{2}(t)^{2}\right)|X^{\xi}(t)-X^{\eta}(t)|^{2}.$
Combining this with (4.10), (4.12), (4.27) and Itô’s formula, for any $k\geq 1$, we find a martingale $M(t)$ and a constant $c_{1}(T,k)>0$ such that for any $t\in[0,T]$,
$\text{\rm{d}}\Xi(t)\leq{c_{1}(T,k)}\left\{\Xi(t)\text{\rm{d}}{A}(t)+\sup_{s\in[0,t]}\Xi(s)\text{\rm{d}}t+\|Y_{0}^{\xi}-Y_{0}^{\eta}\|^{2k}_{\tau}\text{\rm{d}}t\right\}+\text{\rm{d}}M(t).$
Using the Itô formula, the stochastic Grönwall inequality in [3, Lemma A.5] and (4.10), there exists a constant $c_{2}(T,k)>0$ such that
$\left(\mathbb{E}\left[\sup_{s\in[0,t]}\left(\text{\rm{e}}^{-{c_{1}(T,k)}{A}(s)}|\Phi(s)|^{2k}\right)^{\frac{2}{3}}\right]\right)^{\frac{3}{2}}\leq c_{2}(T,k)\|\xi-\eta\|_{\tau}^{2k},\ \ t\in[0,T].$
By combining this with (4.29) and Hölder’s inequality, we find a constant $c_{3}(\varepsilon,T,k)>0$, such that
$\mathbb{E}\left[\sup_{s\in[0,t]}|\Phi(s)|^{k}\right]\leq\mathbb{E}\left[\text{\rm{e}}^{\frac{1}{2}{c_{1}(T,k)}A(t)}\left(\sup_{s\in[0,t]}\text{\rm{e}}^{-\frac{1}{2}{c_{1}(T,k)}A(s)}|\Phi(s)|^{k}\right)\right]$
$\leq\left(\mathbb{E}\left[\text{\rm{e}}^{2{c_{1}(T,k)}A(t)}\right]\right)^{\frac{1}{4}}\left(\mathbb{E}\left[\sup_{s\in[0,t]}\left(\text{\rm{e}}^{-{c_{1}(T,k)}{A}(s)}|\Phi(s)|^{2k}\right)^{\frac{2}{3}}\right]\right)^{\frac{3}{4}}$
$\leq c_{3}(\varepsilon,T,k)\text{\rm{e}}^{\varepsilon(\|\xi\|_{\tau}^{2\alpha}+\|\eta\|_{\tau}^{2\alpha})}\|\xi-\eta\|^{k}_{\tau},\ \ t\in[0,T].$
This together with (4.12) yields the desired estimate (2.4) for some constant $c>0$ depending on $\varepsilon,T$ and $k$. ∎
5 Proof of Theorem 2.3
It suffices to prove the assertions up to any fixed time $T\in(0,\infty).$ To this end, we use the fixed point argument as in [7].
Let $X_{0}$ be an $\mathscr{F}_{0}$-measurable random variable on $\mathscr{C}_{\tau}$ with $\gamma:=\mathscr{L}_{X_{0}}\in\mathscr{P}_{k,\text{\rm{e}}}^{\alpha}(\mathscr{C}_{\tau}).$ For any constant $N\geq 2$, let
$\mathcal{C}^{\gamma,N}_{k}:=\left\{\mu\in C_{b}^{w}([0,T];\mathscr{P}_{k,\text{\rm{e}}}^{\alpha}{(\mathscr{C}_{\tau})}):\mu_{0}=\gamma,\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}(1+\mu_{t}(\|\cdot\|_{\tau}^{k}))\leq N\right\},$
where
$\begin{split}&C^{w}([0,T];\mathscr{P}_{k,\text{\rm{e}}}^{\alpha}(\mathscr{C}_{\tau})):=\{\mu:[0,T]\rightarrow\mathscr{P}_{k,\text{\rm{e}}}^{\alpha}(\mathscr{C}_{\tau}))\text{ is weakly continuous}\},\\ &C^{w}_{b}([0,T];\mathscr{P}_{k,\text{\rm{e}}}^{\alpha}(\mathscr{C}_{\tau})):=\Big{\{}\mu\in C^{w}([0,T];\mathscr{P}_{k,\text{\rm{e}}}^{\alpha}(\mathscr{C}_{\tau})):\sup_{t\in[0,T]}\mathbb{W}_{k,var}(\mu_{t},\mu_{0})<\infty\Big{\}}.\end{split}$
Then as $N\uparrow\infty$,
$\mathcal{C}^{\gamma,N}_{k}\uparrow\mathcal{C}^{\gamma}_{k}:=\left\{\mu\in C_{b}^{w}([0,T];\mathscr{P}_{k,\text{\rm{e}}}^{\alpha}{(\mathscr{C}_{\tau})}):\mu_{0}=\gamma\right\}.$ (5.1)
According to Theorem 2.1 and Corollary 4.5, under the assumptions $(H_{1})$$(H_{2})$, for any $\mu\in\mathcal{C}_{k}^{\gamma}$, the SDE
$\text{\rm{d}}X^{\mu}(t)=b(t,X_{t}^{\mu},\mu_{t})\text{\rm{d}}t+\sigma(t,X^{\mu}(t))\text{\rm{d}}W(t),\ \ t\in[0,T],X_{0}^{\mu}=X_{0}$ (5.2)
has a unique segment solution with
$\Phi_{\cdot}^{\gamma}\mu:=\mathscr{L}_{X_{\cdot}^{\mu}}\in\mathcal{C}_{k}^{\gamma}.$
Then the well-posedness of (1.2) follows if the map $\Phi^{\gamma}$ has a unique fixed point in $\mathcal{C}^{\gamma}_{k}$. To this end, we need to verify that there exists a constant $N_{0}\geq 2$ such that for any $N\geq N_{0},$ the following two assertions hold:
  1. (a)
    $\Phi^{\gamma}:\mathcal{C}_{k}^{\gamma,N}\rightarrow\mathcal{C}_{k}^{\gamma,N},$ i.e.
    $\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\left(1+\|\Phi^{\gamma}_{t}\mu\|_{k}^{k}\right)\leq N,\ \ \mu\in\mathcal{C}^{\gamma,N}_{k}.$ (5.3)
  2. (b)
    $\Phi^{\gamma}$ has a unique fixed point in $\mathcal{C}_{k}^{\gamma,N}.$
Once these two assertions are confirmed, $\Phi^{\gamma}$ has a unique fixed point $\mu\in\mathcal{C}_{k}^{\gamma}$, and $X_{t}=X_{t}^{\mu}$ is the unique segment solution of (1.2) with initial value $X_{0}$, so that (2.7) follows from (5.8).
In the following two subsections, we prove assertions (a) and (b) respectively.
5.1 Proof of (a)
It suffices to prove (5.3) for $k>0$ and large $N$. By Lemma 4.4, $(H_{1})$-$(H_{2})$ implies that for some constant $c_{1}>0$ we have
$\mathbb{E}\left[\exp\left(\int_{0}^{T}|f_{0}(s,X^{\mu}(s))|^{2}\,\text{\rm{d}}s\right)\right]\leq c_{1}\mathbb{E}[\text{\rm{e}}^{\varepsilon\|X_{0}^{\mu}\|_{\tau}^{2\alpha}}]=c_{1}\gamma(\text{\rm{e}}^{\varepsilon\|\cdot\|_{\tau}^{2\alpha}})<\infty.$ (5.4)
Hence,
$\sup_{\mu\in\mathcal{C}_{k}^{\gamma}}\mathbb{E}\bigg{(}\int_{0}^{T}|f_{0}(s,X^{\mu}(s))|^{2}\,\text{\rm{d}}s\bigg{)}^{k}<\infty.$
Combining this with $(H_{1})$-$(H_{2})$, $\gamma:=\mathscr{L}_{X_{0}}\in\mathscr{P}_{k,\text{\rm{e}}}^{\alpha}(\mathscr{C}_{\tau})$, the Itô formula and the BDG inequality, we find constants $c_{2},c_{3}>0$ such that
$\mathbb{E}\left[\sup_{s\in[0,t]}\left(1+\text{\rm{e}}^{k\tau s}|X^{\mu}(s)|^{k}\right)\right]\leq c_{2}\mathbb{E}(1+|X(0)|^{k})$ (5.5)
$+c_{2}\mathbb{E}\left(\int_{0}^{t}\text{\rm{e}}^{\tau s}\left\{\|X^{\mu}_{s}\|_{\tau}+|f_{0}(s,X^{\mu}(s))|+\|\mu_{s}\|_{k}\right\}\,\text{\rm{d}}s\right)^{k}$
$\leq c_{3}+c_{3}\mathbb{E}\left(\int_{0}^{t}\left\{\text{\rm{e}}^{2\tau s}\|X_{s}^{\mu}\|_{\tau}^{2}+\text{\rm{e}}^{2\tau s}\|\mu_{s}\|_{k}^{2}\right\}\text{\rm{d}}s\right)^{\frac{k}{2}},\ \ t\in[0,T].$
Below we verify (5.3) by considering $k\geq 2$ and $k\in(0,2)$ respectively.
$(a_{1})$
Let $k\geq 2$. By (4.12) and (5.5), there exists a constant $c_{4}>0$ such that for any $t\in[0,T]$,
$\mathbb{E}\left[\sup_{s\in[0,t]}\left(1+\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\right)\right]\leq c_{4}+c_{4}\int_{0}^{t}\left\{\mathbb{E}\left[\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\right]+\text{\rm{e}}^{k\tau s}\|\mu_{s}\|_{k}^{k}\right\}\,\text{\rm{d}}s.$
Using Grönwall’s lemma, it holds
$\mathbb{E}\left[\sup_{s\in[0,t]}\left(1+\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\right)\right]\leq c_{4}\text{\rm{e}}^{c_{4}T}\left(1+\int_{0}^{t}\text{\rm{e}}^{k\tau s}\|\mu_{s}\|_{k}^{k}\,\text{\rm{d}}s\right),\ \ t\in[0,T].$
Hence, by noting that $\mu\in\mathcal{C}^{\gamma,N}_{k}$, we can find a constant $c_{5}$ such that
$\begin{split}&\mathbb{E}\left[1+\|X_{t}^{\mu}\|_{\tau}^{k}\right]\leq 1+\text{\rm{e}}^{-k\tau t}\mathbb{E}\left[\sup_{s\in[0,t]}\left(1+\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\right)\right]\\ &\leq c_{5}+c_{5}\int_{0}^{t}\text{\rm{e}}^{-k\tau(t-s)}\|\mu_{s}\|_{k}^{k}\,\text{\rm{d}}s\\ &\leq c_{5}+c_{5}N\text{\rm{e}}^{Nt}\int_{0}^{t}\text{\rm{e}}^{-N(t-s)}\text{\rm{d}}s\leq 2c_{5}\text{\rm{e}}^{Nt},\ \ t\in[0,T].\end{split}$ (5.6)
Taking $N_{0}=2c_{5}$, we derive
$\begin{split}&\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\left(1+\|\Phi_{t}^{\gamma}\mu\|_{k}^{k}\right)=\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\mathbb{E}\left(1+\|X_{t}^{\mu}\|_{\tau}^{k}\right){\leq N},\quad N\geq N_{0},\ \mu\in\mathcal{C}^{\gamma,N}_{k}.\end{split}$
$(a_{2})$
Let $k\in(0,2)$. By (4.12) and (5.5), we find constants $c_{6},c_{7},c_{8}>0$ such that
$\begin{split}&U_{t}:=\mathbb{E}\left[\sup_{s\in[0,t]}\left(1+\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\right)\right]\leq c_{6}+c_{6}\mathbb{E}\left(\int_{0}^{t}\left\{\text{\rm{e}}^{2\tau s}\|X_{s}^{\mu}\|_{\tau}^{2}+\text{\rm{e}}^{2\tau s}\|\mu_{s}\|_{k}^{2}\right\}\,\text{\rm{d}}s\right)^{\frac{k}{2}},\\ &\leq c_{7}+c_{7}\left(\int_{0}^{t}\text{\rm{e}}^{2\tau s}\|\mu_{s}\|_{k}^{2}\,\text{\rm{d}}s\right)^{\frac{k}{2}}+c_{7}\mathbb{E}\left\{\left[\sup_{s\in[0,t]}\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\right]^{1-\frac{k}{2}}\left(\int_{0}^{t}\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\,\text{\rm{d}}s\right)^{\frac{k}{2}}\right\}\\ &\leq c_{7}+c_{7}\left(\int_{0}^{t}\text{\rm{e}}^{2\tau s}\|\mu_{s}\|_{k}^{2}\,\text{\rm{d}}s\right)^{\frac{k}{2}}+\frac{1}{2}U_{t}+c_{8}\int_{0}^{t}U_{s}\,\text{\rm{d}}s,\ \ t\in[0,T],\end{split}$
then Grönwall’s lemma implies that
$U_{t}\leq 2c_{7}\text{\rm{e}}^{2c_{8}T}\left\{1+\left(\int_{0}^{t}\text{\rm{e}}^{2\tau s}\|\mu_{s}\|_{k}^{2}\,\text{\rm{d}}s\right)^{\frac{k}{2}}\right\},\ \ t\in[0,T].$
Thus, there exist constants $c_{9},c_{10}>0$ such that for any $\mu\in\mathcal{C}^{\gamma,N}_{k}$,
$\begin{split}&\mathbb{E}\left(1+\|X_{t}^{\mu}\|_{\tau}^{k}\right)\leq 1+\text{\rm{e}}^{-k\tau t}U_{t}\leq c_{9}+c_{9}\left(\int_{0}^{t}\text{\rm{e}}^{-2\tau(t-s)}\|\mu_{s}\|_{k}^{2}\,\text{\rm{d}}s\right)^{\frac{k}{2}}\\ &\leq c_{9}+c_{9}Ne^{Nt}\left(\int_{0}^{t}\text{\rm{e}}^{-2N(t-s)}\,\text{\rm{d}}s\right)^{\frac{k}{2}}\leq c_{9}+c_{10}N^{1-\frac{k}{2}}\text{\rm{e}}^{Nt},\ \ t\in[0,T].\end{split}$ (5.7)
Thus, there exists a constant $N_{0}>2$ such that for any $N\geq N_{0}$,
$\begin{split}&\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\left(1+\|\Phi_{t}^{\gamma}\mu\|_{k}^{k}\right)=\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\mathbb{E}\left(1+\|X_{t}^{\mu}\|_{\tau}^{k}\right)\\ &\leq c_{9}+c_{10}N^{1-\frac{k}{2}}\leq N,\ \ \mu\in\mathcal{C}^{\gamma,N}_{k}.\end{split}$
5.2 Proof of (b)
To ensure that $\Phi^{\gamma}$ has a unique fixed point in $\mathcal{C}_{k}^{\gamma,N}$ for $N\geq N_{0}$, we shall prove that it is contractive under a complete metric.
For any $\theta>0$, let
$\mathbb{W}_{k,\theta,var}(\mu,\nu):=\sup_{t\in[0,T]}\text{\rm{e}}^{-\theta t}\|\mu_{t}-\nu_{t}\|_{k,var},$
$\mathbb{W}_{k,\theta}(\mu,\nu):=\sup_{t\in[0,T]}\text{\rm{e}}^{-\theta t}\mathbb{W}_{k}(\mu_{t},\nu_{t}),\ \ \mu,\nu\in\mathcal{C}^{\gamma,N}_{k}.$
Then the metric $\tilde{\mathbb{W}}_{k,\theta,var}=\mathbb{W}_{k,\theta,var}+\mathbb{W}_{k,\theta}$ is complete on $\mathcal{C}_{k}^{\gamma,N}$. We intend to show that $\Phi^{\gamma}$ is contractive in $C^{\gamma,N}_{k}$ under the metric $\tilde{\mathbb{W}}_{k,\theta,var}$ when $\theta$ is large enough.
By Theorem 2.1, $(H_{1})$-$(H_{2})$, (5.3) and $\gamma:=\mathscr{L}_{X_{0}}\in\mathscr{P}_{k}(\mathscr{C}_{\tau})$, we find a constant $C_{0}(N)>0$ such that
$\sup_{\mu\in C_{k}^{\gamma,N}}\mathbb{E}\Big{(}\mathbb{E}\Big{[}\sup_{t\in[0,T]}\|X^{\mu}_{t}\|_{\tau}^{2k}\Big{|}\mathscr{F}_{0}\Big{]}\Big{)}^{\frac{1}{2}}\leq C_{0}(N)\mathbb{E}\left(1+\|X_{0}\|_{\tau}^{k}\right)<\infty.$ (5.8)
Next, since $\mu,\nu\in\mathcal{C}^{\gamma,N}_{k}$, we find a constant $C_{0}(N)>0$ such that
$\sup_{t\in[0,T]}\mathbb{W}_{k,var}(\mu_{t},\nu_{t})\leq C_{0}(N).$
Thus, by $(H_{1})$ and $(H_{2})$, we find a constant $C_{1}(N)>0$ such that for any $\mu,\nu\in\mathcal{C}^{\gamma,N}_{k}$,
$\zeta_{s}:=\left\{\sigma^{*}(\sigma\sigma^{*})^{-1}\right\}(s,X^{\mu}(s))\left[b^{(1)}(s,X_{s}^{\mu},\mu_{s})-b^{(1)}(s,X_{s}^{\mu},\nu_{s})\right],\ \ s\in[0,T]$
satisfies
$|\zeta_{s}|^{2}\leq C_{1}(N)H(s)\left(1\wedge\mathbb{W}_{k,var}(\mu_{s},\nu_{s})^{2}\right),\ \ s\in[0,T].$ (5.9)
Recalling that $H\in L_{loc}^{1}([0,\infty);(0,\infty))$, we have
$\int_{0}^{t}H(s)\left(1\wedge\mathbb{W}_{k,var}(\mu_{s},\nu_{s})^{2}\right)\,\text{\rm{d}}s\leq\int_{0}^{t}H(s)\,\text{\rm{d}}s<\infty,$
then by $(H_{2})$ and Girsanov’s theorem, for any $t\in[0,T]$,
$R_{t}:=\exp\left(\int_{0}^{t}\langle\zeta_{s},\text{\rm{d}}W(s)\rangle-\frac{1}{2}\int_{0}^{t}|\zeta_{s}|^{2}\,\text{\rm{d}}s\right)$ (5.10)
is a martingale and
$\tilde{W}_{u}:=W_{u}-\int_{0}^{u}\zeta_{s}\,\text{\rm{d}}s,\ \ u\in[0,t]$
is a Brownian motion under the probability measure $\mathbb{Q}_{t}:=R(t)\mathbb{P}$. Since $\text{\rm{e}}^{s}-1\leq s\text{\rm{e}}^{s}$ for $s\geq 0$, we find constants $C_{2}(N),C_{3}(N)>0$ such that
$\begin{split}&\mathbb{E}\left(|R_{t}-1|^{2}\big{|}\mathscr{F}_{0}\right)=\mathbb{E}\left[\text{\rm{e}}^{2\int_{0}^{t}\langle\zeta_{s},\text{\rm{d}}W(s)\rangle-2\int_{0}^{t}|\zeta_{s}|^{2}\,\text{\rm{d}}s+\int_{0}^{t}|\zeta_{s}|^{2}\,\text{\rm{d}}s}\bigg{|}\mathscr{F}_{0}\right]-1\\ &\leq\text{\rm{e}}^{C_{1}(N)\int_{0}^{t}H(s)\left(1\wedge\mathbb{W}_{k,var}(\mu_{s},\nu_{s})^{2}\right)\,\text{\rm{d}}s}-1\leq C_{2}(N)\int_{0}^{t}H(s)\mathbb{W}_{k,var}(\mu_{s},\nu_{s})^{2}\,\text{\rm{d}}s\\ &\leq C_{3}(N)\text{\rm{e}}^{2\theta t}\tilde{\mathbb{W}}_{k,\theta,var}(\mu,\nu)^{2}\int_{0}^{t}H(s)\text{\rm{e}}^{-2\theta(t-s)}\,\text{\rm{d}}s,\ \ \mu,\nu\in\mathcal{C}^{\gamma,N}_{k}.\end{split}$ (5.11)
Reformulating (5.2) as
$\text{\rm{d}}X^{\mu}(r)=b(r,X_{r}^{\mu},\nu_{r})\text{\rm{d}}r+\sigma(r,X^{\mu}(r))\text{\rm{d}}\tilde{W}(r),\ \ \mathscr{L}_{X_{0}^{\mu}}=\gamma,r\in[0,t].$ (5.12)
By the uniqueness, we obtain
$\Phi_{t}^{\gamma}\nu=\mathscr{L}_{X_{t}^{\nu}}=\mathscr{L}_{X_{t}^{\mu}|\mathbb{Q}_{t}},$
where $\mathscr{L}_{X_{t}^{\mu}|\mathbb{Q}_{t}}$ stands for the distribution of $X_{t}^{\mu}$ under $\mathbb{Q}_{t}$. Thus, by (5.8), (5.9), (LABEL:N1) and Hölder’s inequality, we find constant $C_{4}(N)>0$ such that
$\|\Phi^{\gamma}_{t}\mu-\Phi^{\gamma}_{t}\nu\|_{k,var}=\sup_{|f|\leq 1+\|\cdot\|_{\tau}^{k}}\left|\mathbb{E}\left[f(X_{t}^{\nu})-f(X_{t}^{\mu})\right]\right|$
$=\sup_{|f|\leq 1+\|\cdot\|_{\tau}^{k}}\left|\mathbb{E}\left[(R_{t}-1)f(X_{t}^{\mu})\right]\right|\leq\mathbb{E}\left[(1+\|X_{t}^{\mu}\|_{\tau}^{k})|R_{t}-1|\right]$
$\leq\mathbb{E}\left[\left\{\mathbb{E}\left((1+\|X_{t}^{\mu}\|_{\tau}^{k})^{2}\big{|}\mathscr{F}_{0}\right)\right\}^{\frac{1}{2}}\left\{\mathbb{E}\left(|R_{t}-1|^{2}\big{|}\mathscr{F}_{0}\right)\right\}^{\frac{1}{2}}\right]$
$\leq C_{4}(N)\text{\rm{e}}^{\theta t}\tilde{\mathbb{W}}_{k,\theta,var}(\mu,\nu)\left(\int_{0}^{t}H(s)\text{\rm{e}}^{-2\theta(t-s)}\,\text{\rm{d}}s\right)^{\frac{1}{2}},\ \ \mu,\nu\in\mathcal{C}^{\gamma,N}_{k}.$
Therefore,
$\mathbb{W}_{k,\theta,var}(\Phi^{\gamma}\mu,\Phi^{\gamma}\nu)=\sup_{t\in[0,T]}\text{\rm{e}}^{-\theta t}\|\Phi^{\gamma}_{t}\mu-\Phi^{\gamma}_{t}\nu\|_{k,var}$ (5.13)
$\leq C_{4}(N)\sup_{t\in[0,T]}\left(\int_{0}^{t}H(s)\text{\rm{e}}^{-2\theta(t-s)}\,\text{\rm{d}}s\right)^{\frac{1}{2}}\tilde{\mathbb{W}}_{k,\theta,var}(\mu,\nu),\ \ \mu,\nu\in\mathcal{C}^{\gamma,N}_{k}.$
We will finish the proof by considering $k\leq 1$ and $k>1$ respectively.
$(b_{1})$
Let $k\leq 1$. By (1.3) and (5.13), we obtain
$\tilde{\mathbb{W}}_{k,\theta,var}(\Phi^{\gamma}\mu,\Phi^{\gamma}\nu)\leq(1+c)\mathbb{W}_{k,\theta,var}(\Phi^{\gamma}\mu,\Phi^{\gamma}\nu)$
$\leq(1+c)C_{4}(N)\left(\int_{0}^{t}H(s)\text{\rm{e}}^{-2\theta(t-s)}\,\text{\rm{d}}s\right)^{\frac{1}{2}}\tilde{\mathbb{W}}_{k,\theta,var}(\mu,\nu),\ \ \mu,\nu\in\mathcal{C}^{\gamma,N}_{k}.$
Noting that $H\in L_{loc}^{1}([0,\infty);(0,\infty))$ yields that
$\lim_{\theta\rightarrow\infty}\sup_{t\in[0,T]}\left(\int_{0}^{t}H(s)\text{\rm{e}}^{-2\theta(t-s)}\,\text{\rm{d}}s\right)^{\frac{1}{2}}=0,$ (5.14)
we may choose $\theta>0$ large enough such that
$\tilde{\mathbb{W}}_{k,\theta,var}(\Phi^{\gamma}\mu,\Phi^{\gamma}\nu)\leq\frac{1}{2}\tilde{\mathbb{W}}_{k,\theta,var}(\mu,\nu),\ \ \mu,\nu\in\mathcal{C}^{\gamma,N}_{k}.$
This together with (a) implies that $\Phi^{\gamma}$ has a unique fixed point in $C^{\gamma,N}_{k}.$
$(b_{2})$
Let $k>1$. Let $\Theta(t,\cdot):=id+u(t,\cdot)$ for $u$ solving (4.7) such that (4.8) and (4.9) hold. Let
$Z(t):={X^{\mu}(t)+\Theta(t,X^{\mu}(t))-X^{\nu}(t)-\Theta(t,X^{\nu}(t))},$
$g_{1}(t):=\mathscr{M}\|\nabla^{2}u\|(t,X^{\mu}(t))+\mathscr{M}\|\nabla^{2}u\|(t,X^{\nu}(t)),$
$g_{2}(t):=g_{1}(t)+\mathscr{M}\|\nabla\sigma\|(t,X^{\mu}(t))+\mathscr{M}\|\nabla\sigma\|(t,X^{\nu}(t)),\ \ \ t\in[0,T].$
By $(H_{1})$-$(H_{3})$, (4.9), (4.10) and Itô’s formula, we find a constant $c_{1}(k)>0$ such that
$\text{\rm{d}}\big{\{}\text{\rm{e}}^{2k\tau t}|Z(t)|^{2k}\big{\}}\leq c_{1}(k)\left\{\text{\rm{e}}^{2k\tau t}\|Z_{t}\|_{\tau}^{2k}+H(t)\text{\rm{e}}^{2k\tau t}\mathbb{W}_{k,var}(\mu_{t},\nu_{t})^{2k}\right\}\text{\rm{d}}t$ (5.15)
$+\text{\rm{e}}^{2k\tau t}|Z(t)|^{2k}\text{\rm{d}}A(t)+\text{\rm{d}}M(t),\ \ t\in[0,T],$
where
$\begin{split}&A(t):={c_{1}(k)}\int_{0}^{t}\left(1+g_{2}(s)^{2}+g_{1}(s)\|X^{\mu}_{s}\|_{\tau}^{\alpha}\right)\,\text{\rm{d}}s,\\ &\text{\rm{d}}\langle M\rangle(t)\leq c_{1}(k)\left(1+g_{2}(t)\right)^{2}\text{\rm{e}}^{4k\tau t}|Z(t)|^{4k}\,\text{\rm{d}}t.\end{split}$ (5.16)
Combining this with (4.10), (4.12), the stochastic Grönwall inequality in [3, Lemma A.5] , Lemma 4.4, Corollary 4.5 and the fact that $X_{0}^{\mu}=X_{0}^{\nu}=X_{0}$, we find a constant $c_{2}(k)>0$ such that
$\begin{split}&\mathbb{E}\left[\sup_{s\in[0,t]}\text{\rm{e}}^{k\tau s}\|X^{\mu}_{s}-X^{\nu}_{s}\|_{\tau}^{k}\right]{=\mathbb{E}\left[\sup_{s\in[0,t]}\text{\rm{e}}^{k\tau s}|X^{\mu}(s)-X^{\nu}(s)|^{k}\right]}\\ &\leq\left(\mathbb{E}\left[\text{\rm{e}}^{\frac{3}{2}A(t)}\right]\right)^{\frac{1}{3}}\left(\mathbb{E}\left[\sup_{s\in[0,t]}\text{\rm{e}}^{-A(s)}\text{\rm{e}}^{2k\tau s}{|X^{\mu}(s)-X^{\nu}(s)|}^{2k}\right]^{\frac{3}{4}}\right)^{\frac{2}{3}}\\ &\leq c_{2}(k)\left(\int_{0}^{t}H(s)\text{\rm{e}}^{2k\tau s}\mathbb{W}_{k,var}(\mu_{s},\nu_{s})^{2k}\,\text{\rm{d}}s\right)^{\frac{1}{2}}.\end{split}$
Since $\Phi_{t}^{\gamma}\mu=\mathscr{L}_{X^{\mu}_{t}}$ and $\Phi_{t}^{\gamma}\nu=\mathscr{L}_{X^{\nu}_{t}},$ by the inequalities above, we find a constant $c_{3}(k)>0$ such that
$\mathbb{W}_{k,\theta}(\Phi^{\gamma}\mu,\Phi^{\gamma}\nu)=\sup_{t\in[0,T]}\text{\rm{e}}^{-\theta t}\mathbb{W}_{k}(\Phi_{t}^{\gamma}\mu,\Phi_{t}^{\gamma}\nu)\leq\sup_{t\in[0,T]}\text{\rm{e}}^{-\theta t}\left(\mathbb{E}\left[\|X_{t}^{\mu}-X_{t}^{\nu}\|_{\tau}^{k}\right]\right)^{\frac{1}{k}}$
$\leq c_{2}(k)\sup_{t\in[0,T]}\left(\int_{0}^{t}H(s)\text{\rm{e}}^{-2k\theta t}\mathbb{W}_{k,var}(\mu_{s},\nu_{s})^{2k}\,\text{\rm{d}}s\right)^{\frac{1}{2k}}$
$\leq c_{3}(k)\tilde{\mathbb{W}}_{k,\theta,var}(\mu,\nu)\sup_{t\in[0,T]}\left(\int_{0}^{t}H(s)\text{\rm{e}}^{-2k\theta(t-s)}\,\text{\rm{d}}s\right)^{\frac{1}{2k}},\ \ \mu,\nu\in\mathcal{C}^{\gamma,N}_{k},\ \theta>0.$
Combining this with (5.13) and (5.14), we may choose $\theta>0$ such that
$\tilde{\mathbb{W}}_{k,\theta,var}(\Phi^{\gamma}\mu,\Phi^{\gamma}\nu)\leq\frac{1}{2}\tilde{\mathbb{W}}_{k,\theta,var}(\mu,\nu),\ \ \mu,\nu\in\mathcal{C}^{\gamma,N}_{k}.$
Thus, $\Phi^{\gamma}$ has a unique fixed point in $\mathcal{C}^{\gamma,N}_{k}$.
6 Proof of Theorem 3.3
Proof of Theorem 3.3(1).
The well-posedness can be proved by using a standard fixed point theorem. Let $X_{0}$ be $\mathscr{F}_{0}$-measurable with $\gamma:=\mathscr{L}_{X_{0}}\in\mathscr{P}_{2}(\mathscr{C}_{\tau})$. For any constant $\theta,T>0,$ the path space
$\mathcal{C}_{2}^{\gamma}:=\big{\{}\mu\in C^{w}([0,T];\mathscr{P}_{2}(\mathscr{C}_{\tau})):\ \mu_{0}=\gamma\big{\}}$
is complete under the metric
$\mathbb{W}_{2,\theta}(\mu,\nu):=\sup_{t\in[0,T]}\text{\rm{e}}^{-\theta t}\mathbb{W}_{2}(\mu_{t},\nu_{t}).$
According to Theorem 3.1, $({H}^{\prime})$ implies that for any $\mu\in\mathcal{C}_{2}^{\gamma}$, the SDE
$\text{\rm{d}}X^{\mu}(t)=b(t,X_{t}^{\mu},\mu_{t})\text{\rm{d}}t+\sigma(t,X^{\mu}_{t})\text{\rm{d}}W(t),\ \ t\in[0,T],X_{0}^{\mu}=X_{0}$ (6.1)
is well-posed with
$\Phi_{\cdot}^{\gamma}\mu:=\mathscr{L}_{X_{\cdot}^{\mu}}\in\mathcal{C}_{2}^{\gamma}.$
For the well-posedness of (1.2), it suffices to show that $\Phi^{\gamma}$ has a unique fixed point in $\mathcal{C}_{2}^{\gamma}$.
Now, for $\nu_{0}\in\mathscr{P}_{2}(\mathscr{C}_{\tau})$, we choose $\mathscr{F}_{0}$-measurable random variables $X_{0}^{\mu}$ and $X_{0}^{\nu}$ on $\mathscr{C}_{\tau}$ such that
$\mathbb{W}_{2}(\mu_{0},\nu_{0})^{2}=\mathbb{E}\big{[}\|X_{0}^{\mu}-X_{0}^{\nu}\|_{\tau}^{2}\big{]}.$ (6.2)
For any $\nu\in\mathcal{C}_{2}^{\nu_{0}}$ which is defined as $\mathcal{C}_{2}^{\gamma}$ for $\nu_{0}$ replacing $\gamma$, let $X_{t}^{\nu}$ be the unique solution to
$\text{\rm{d}}X^{\nu}(t)=b(t,X_{t}^{\nu},\nu_{t})\text{\rm{d}}t+\sigma(t,X^{\nu}_{t})\text{\rm{d}}W(t),\ \ t\in[0,T]$ (6.3)
with initial value $X_{0}^{\nu}$. By $(H^{\prime})$ and Itô’s formula, we obtain
$\text{\rm{d}}{\text{\rm{e}}^{2\tau t}}|X^{\mu}(t)-X^{\nu}(t)|^{2}\leq K{\text{\rm{e}}^{2\tau t}}\big{(}\|X_{t}^{\mu}-X_{t}^{\nu}\|_{\tau}^{2}+\mathbb{W}_{2}(\mu_{t},\nu_{t})^{2}\big{)}\text{\rm{d}}t+\text{\rm{d}}M(t)$
for some constant $K>0$, where $M_{t}$ is a martingale with
$\text{\rm{d}}\langle M\rangle(t)\leq K{\text{\rm{e}}^{4\tau t}}\|X^{\mu}_{t}-X_{t}^{\nu}\|_{\tau}^{4}\text{\rm{d}}t.$
Combining this with (4.12), Itô’s isometry and Young’s inequality, we find a constant $c_{1}>0$ such that
$\mathbb{E}\bigg{[}\sup_{s\in[0,t]}{\text{\rm{e}}^{2\tau s}}\|X_{s}^{\mu}-X_{s}^{\nu}\|_{\tau}^{2}\bigg{]}\leq c_{1}\mathbb{E}\|X_{0}^{\mu}-X_{0}^{\nu}\|_{\tau}^{2}+c_{1}\mathbb{E}\bigg{[}\bigg{(}\int_{0}^{t}{\text{\rm{e}}^{4\tau s}}\|X_{s}^{\mu}-X_{s}^{\nu}\|_{\tau}^{4}\,\text{\rm{d}}s\bigg{)}^{\frac{1}{2}}\bigg{]}$
$\qquad\qquad\qquad\qquad\qquad\qquad+c_{1}\int_{0}^{t}{\text{\rm{e}}^{2\tau s}}\mathbb{E}\big{[}\|X_{s}^{\mu}-X_{s}^{\nu}\|_{\tau}^{2}+\mathbb{W}_{2}(\mu_{s},\nu_{s})^{2}\big{]}\,\text{\rm{d}}s$
$\leq c_{1}\mathbb{E}\|X_{0}^{\mu}-X_{0}^{\nu}\|_{\tau}^{2}+\Big{(}c_{1}+\frac{c_{1}^{2}}{2}\Big{)}\int_{0}^{t}{\text{\rm{e}}^{2\tau s}}\mathbb{E}\big{[}\|X_{s}^{\mu}-X_{s}^{\nu}\|_{\tau}^{2}+\mathbb{W}_{2}(\mu_{s},\nu_{s})^{2}\big{]}\,\text{\rm{d}}s$
$\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{2}\mathbb{E}\bigg{[}\sup_{s\in[0,t]}{\text{\rm{e}}^{2\tau s}}\|X_{s}^{\mu}-X_{s}^{\nu}\|_{\tau}^{2}\bigg{]},\ \ t\in[0,T].$
By an approximation argument with stopping times, we may and do assume that
$\mathbb{E}\bigg{[}\sup_{s\in[0,t]}{\text{\rm{e}}^{2\tau s}}\|X_{s}^{\mu}-X_{s}^{\nu}\|_{\tau}^{2}\bigg{]}<\infty,$
so that the above estimate together with (6.2) and Grönwall’s inequality yields
$\mathbb{E}\bigg{[}\sup_{s\in[0,t]}\|X_{s}^{\mu}-X_{s}^{\nu}\|_{\tau}^{2}\bigg{]}\leq c_{2}(T)\mathbb{W}_{2}(\mu_{0},\nu_{0})^{2}+c_{2}(T)\int_{0}^{t}\mathbb{W}_{2}(\mu_{s},\nu_{s})^{2}\,\text{\rm{d}}s$ (6.4)
for any $t\in[0,T]$ and some constant $c_{2}(T)>0$. In particular, when $\mu_{0}=\nu_{0}=\gamma$, we derive
$\mathbb{W}_{2,\theta}(\Phi^{\gamma}\mu,\Phi^{\gamma}\nu)^{2}\leq\sup_{t\in[0,T]}\text{\rm{e}}^{-\theta t}\mathbb{E}\big{[}\|X_{t}^{\mu}-X_{t}^{\nu}\|_{\tau}^{2}\big{]}$
$\leq c_{2}(T)\sup_{t\in[0,T]}\text{\rm{e}}^{-\theta t}\int_{0}^{t}\mathbb{W}_{2}(\mu_{s},\nu_{s})^{2}\,\text{\rm{d}}s\leq\frac{c_{2}(T)}{\theta}\mathbb{W}_{2,\theta}(\mu,\nu)^{2},\ \ \mu,\nu\in\mathcal{C}_{2}^{\gamma}.$
So, when $\theta>c_{2}(T)$, $\Phi^{\gamma}$ is contractive in the complete metric space $(\mathcal{C}_{2}^{\gamma},\mathbb{W}_{2,\theta})$, hence it has a unique fixed point.
Moreover, letting $\mu_{t}=P_{t}^{*}\mu,\nu_{t}=P_{t}^{*}\nu$, we have $\mathscr{L}_{X_{t}^{\mu}}=P_{t}^{*}\mu$ and $\mathscr{L}_{X_{t}^{\nu}}=P_{t}^{*}\nu$, so that (6.4) and Grönwall’s inequality yield (3.4) for some constant $c>0$. ∎
Now, let $\tau_{0}\in(0,\tau)$ and $\|\sigma\|_{\infty}+\|\sigma^{-1}\|_{\infty}<\infty,$ it remains to verify (3.5), which implies (3.6) through repeating the proof of [5, Theorem 2.1(1)] for
$E=\mathscr{P}_{2,\text{\rm{e}}}^{\alpha}(\mathscr{C}_{\tau}),\ \rho(\mu,\nu)=\mathbb{W}_{2}(\mu,\nu),\ \Gamma_{t}=c\text{\rm{e}}^{-\tau_{0}t},$ (6.5)
and $\Lambda_{t}=c\text{\rm{e}}^{ct}$ in place of $\Lambda$.
Proof of (3.5).
Let $X_{t}^{\mu}$ be the unique solution to (1.2) with the initial distribution $\mu$ and denote $\mu_{t}=P_{t}^{*}\mu$, $\nu_{t}=P_{t}^{*}\nu$,
$\bar{\zeta}_{s}:=\left\{\sigma^{*}(\sigma\sigma^{*})^{-1}\right\}(s,X^{\mu}_{s})\left[b(s,X^{\mu}_{s},\mu_{s})-b(s,X^{\mu}_{s},\nu_{s})\right],$
$\bar{R}_{s}:=\exp\left[-\int_{0}^{s}\langle\bar{\zeta}_{u},\text{\rm{d}}W(u)\rangle-\frac{1}{2}\int_{0}^{s}|\bar{\zeta}_{u}|^{2}\,\text{\rm{d}}u\right],\ \ s\in[0,t].$
Then $(H^{\prime})$ and (3.4) implies that
$|\bar{\zeta}_{s}|^{2}\leq c_{1}K_{2}\mathbb{W}_{2}(\mu_{s},\nu_{s})^{2}\leq c_{2}K_{2}\text{\rm{e}}^{c_{2}s}\mathbb{W}_{2}(\mu,\nu)^{2}$ (6.6)
for some constants $c_{1},c_{2}>0$. Thus by Girsanov’s theorem, $\bar{R}_{s}$ is a martingale and
$\bar{W}(s)=W(s)+\int_{0}^{s}\bar{\zeta}_{r}\,\text{\rm{d}}r,\ \ s\in[0,t]$
is a Brownian motion under the probability measure ¯$\bar{\mathbb{P}}:=\bar{R}_{t}\mathbb{P}$. Then (1.2) can be reformulated as
$\text{\rm{d}}X^{\mu}(s)=b(s,X^{\mu}_{s},\nu_{s})\text{\rm{d}}s+\sigma(s,X^{\mu}_{s})\text{\rm{d}}\bar{W}(s),\ \ \mathscr{L}_{X^{\mu}_{0}}=\mu,\ s\in[0,t].$ (6.7)
For any $\kappa>\tau$, where $\tau>0$ is given in (1.1), consider the following SDE:
$\text{\rm{d}}Y(s)=$$\{b(s,Y_{s},\nu_{s})+\kappa\sigma(s,Y_{s})\sigma(s,X^{\mu}_{s})^{-1}(X^{\mu}(s)-Y(s))\}\text{\rm{d}}s$ (6.8)
$+\sigma(s,Y_{s})\text{\rm{d}}\bar{W}(s),\ \ s\in[0,t],\ \ Y_{0}=X_{0}^{\nu}.$
Let
$\tilde{\zeta}_{s}:=\kappa\sigma(s,X^{\mu}_{s})^{-1}(X^{\mu}(s)-Y(s)),$
$\tilde{R}_{s}:=\exp\left[-\int_{0}^{s}\langle\tilde{\zeta}_{r},\text{\rm{d}}\bar{W}(r)\rangle-\frac{1}{2}\int_{0}^{s}|\tilde{\zeta}_{r}|^{2}\,\text{\rm{d}}r\right],\ \ s\in[0,t].$
Due to [5, Lemma 3.2] with $\mathbb{P}$ replaced by $\bar{\mathbb{P}}$, under the assumption $(H^{\prime})$,
$\tilde{W}(s)=\bar{W}(s)+\int_{0}^{s}\tilde{\zeta}_{r}\,\text{\rm{d}}r=W(s)+\int_{0}^{s}\left(\tilde{\zeta}_{r}+\bar{\zeta}_{r}\right)\,\text{\rm{d}}r,\ \ s\in[0,t]$
is a Brownian motion under the probability measure $\mathbb{Q}=\tilde{R}_{t}\bar{\mathbb{P}}=\tilde{R}_{t}\bar{R}_{t}\mathbb{P}$. Hence (6.8) can be reformulated as
$\text{\rm{d}}Y(s)=b(s,Y_{s},\nu_{s})\text{\rm{d}}s+\sigma(s,Y_{s})\text{\rm{d}}\tilde{W}(s),\ \ s\in[0,t],\ \ Y_{0}=X_{0}^{\nu},$
which together with the uniqueness of (6.3) derives that $\mathscr{L}_{Y_{t}|\mathbb{Q}}=\mathscr{L}_{X_{t}^{\nu}}$. Moreover, if we choose $\mathscr{F}_{0}$-measurable random variables $X_{0}^{\mu}$ and $X_{0}^{\nu}$ on $\mathscr{C}_{\tau}$ such that
$\mathbb{W}_{2}(\mu,\nu)^{2}=\mathbb{E}\big{[}\|X_{0}^{\mu}-X_{0}^{\nu}\|_{\tau}^{2}\big{]}.$ (6.9)
By (6.7), $(H^{\prime})$ and the proof of [5, Lemma 3.3], for any $p>0$ and $\tau_{0}\in(0,\tau)$ we find a constant $\kappa>\tau$ to define $Y_{s}$ in (6.8) such that
$\mathbb{E}_{\mathbb{Q}}\left[\|X^{\mu}_{t}-Y_{t}\|_{\tau}^{p}|\mathscr{F}_{0}\right]\leq c\text{\rm{e}}^{-p\tau_{0}t}\|X^{\mu}_{0}-X^{\nu}_{0}\|_{\tau}^{p},\ \ t\geq 0$ (6.10)
holds for some constant $c>0$. Therefore, applying Young’s inequality in [2, Lemma 2.4],
$\begin{split}&P_{t}\log f(\nu)=\mathbb{E}_{\mathbb{Q}}\left[\log f(Y_{t})\right]=\mathbb{E}_{\mathbb{Q}}\left[\log f(X^{\mu}_{t})\right]+\mathbb{E}_{\mathbb{Q}}\left[\log f(Y_{t})-\log f(X^{\mu}_{t})\right]\\ &\leq\mathbb{E}\left[\bar{R}_{t}\tilde{R}_{t}\log f(X^{\mu}_{t})\right]+\|\nabla\log f\|_{\infty}\mathbb{E}_{\mathbb{Q}}\|X_{t}^{\mu}-Y_{t}\|_{\tau}\\ &\leq\mathbb{E}\left[\bar{R}_{t}\tilde{R}_{t}\log\left(\bar{R}_{t}\tilde{R}_{t}\right)\right]+\log P_{t}f(\mu)+c\|\nabla\log f\|_{\infty}\text{\rm{e}}^{-\tau_{0}t}\mathbb{W}_{2}(\mu,\nu),\ \ t\geq 0\end{split}$ (6.11)
holds for any $f\in\mathscr{B}_{b}^{+}(\mathscr{C}_{r})$ with $\|\nabla\log f\|_{\infty}<\infty$. Next, denote $R_{t}=\bar{R}_{t}\tilde{R}_{t}$, it follows from (6.6), (6.10) that for some positive constants $c_{3},c_{4}$,
$\mathbb{E}\left[R_{t}\log R_{t}\right]\leq\frac{1}{2}\mathbb{E}_{\mathbb{Q}}\int_{0}^{t}|\bar{\zeta}_{s}+\tilde{\zeta}_{s}|^{2}\,\text{\rm{d}}s\leq\mathbb{E}_{\mathbb{Q}}\int_{0}^{t}|\bar{\zeta}_{s}|^{2}\,\text{\rm{d}}s+\mathbb{E}_{\mathbb{Q}}\int_{0}^{t}|\tilde{\zeta}_{s}|^{2}\,\text{\rm{d}}s$
$\leq c_{3}K_{2}\text{\rm{e}}^{c_{3}t}\mathbb{W}_{2}(\mu,\nu)^{2}+c_{3}\int_{0}^{t}\mathbb{E}_{\mathbb{Q}}\|X^{\mu}_{s}-Y_{s}\|_{\tau}^{2}\,\text{\rm{d}}s\leq\left(c_{3}K_{2}\text{\rm{e}}^{c_{3}t}+c_{4}\right)\mathbb{W}_{2}(\mu,\nu)^{2}.$
Substituting this back into (6.11) yields (3.5). ∎
Acknowledgements
The authors would like to thank the associated editor and referees for their helpful comments and suggestions.